### v1

# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

Abstract
We present a fine-grained theoretical analysis of the performance gap between reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) under a representation gap. Our study decomposes this gap into two sources: an explicit representation gap under exact optimization and an implicit representation gap under finite samples. In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities. We show that RLHF, DPO, or online DPO can outperform one another depending on the type of model mis-specifications. Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified. In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model—highlighting a statistical advantage of two-stage learning. Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
[…]
1 Introduction
Reinforcement learning from human feedback (RLHF, Christiano et al. (2017); Ziegler et al. (2019)) is an important paradigm improving the natural language understanding and generation capabilities of large language models (LLMs). The core idea of RLHF is to utilize pair-wise comparison between responses from human annotators, as directly collecting absolute reward signals is hard. There are two stages in RLHF: the reward modeling stage and the policy optimization stage. The reward modeling stage assumes human preferences follow the Bradley-Terry (BT) model (Bradley and Terry, 1952), allowing a prompt-response pair to be assigned a scalar reward. Thus, a reward model $r_{\phi}$ could be trained using negative log-likelihood loss function from human preferences. In the policy optimization stage, the base LM is “online” fine-tuned with RL algorithms such as proximal policy optimization (PPO, Schulman et al. (2017)), based on $r_{\phi}$ under a Kullback-Leibler (KL) divergence-regularized multi-armed bandit setting. And the key assumption behind this two-stage pipeline is the realizability of the ground-truth reward.
The above RLHF paradigm falls inside a broader problem, preference-based policy learning (Wirth et al., 2017). Another popular algorithm in this area is direct preference optimization (DPO, Rafailov et al. (2023)), which utilizes the closed-form solution (assuming realizability as well) for the policy optimization stage to bypass the reward modeling stage and directly fine-tune the base LM as a policy model $\pi_{\theta}$ using the preference dataset. Due to its inherent supervised learning (offline and RL-free) nature, DPO training is more stable than RLHF. And its iterative online version (Guo et al., 2024; Dong et al., 2024) has been shown to have better convergence rates (Shi et al., 2025), and milder coverage conditions (Song et al., 2024; Xiong et al., 2024), than vanilla DPO. The key assumption behind DPO’s design is the realizability of the closed-form solution of the optimal policy.
Notably, in the foundational work of preference learning (Zhu et al., 2023), the ground-truth reward is assumed to lie in a linear model class; and in Rafailov et al. (2023), both the reward class and policy class are tabular parameterized, making their optimal solutions realizable. The realizability condition is commonly assumed in theoretical studies of preference learning (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Yao et al., 2025; Swamy et al., 2025), or DPO-style algorithm designs to derive the loss functions for neural policy classes (Azar et al., 2023; Zhou et al., 2024; Liu et al., 2024b; Xu et al., 2024a). Importantly, under the realizability assumption, it is straightforward to derive the equivalence between the ideal performances of RLHF and DPO (Swamy et al., 2025).
However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., $6$B vs. $175$B in Ouyang et al. (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily mis-specified due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data. Consequently, one should not expect DPO to perform identically to RLHF. This motivates the central question of our investigation:
Under what conditions is DPO equivalent, superior, or inferior to RLHF in performance?
To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($y$ is a response and we omit prompts): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{y\sim\pi}[r^{\star}(y)]-\beta\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)$, where $\pi$ is a distribution over responses, $\pi_{{\mathsf{ref}}}$ is the base model. Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
Our contributions. We study the performance differences between two-stage RLHF and DPO under a representation gap, from an optimization perspective. Our contributions are listed as follows:
Sec. 3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi)\propto-V_{r^{\star}}^{\pi}+\mathcal{O}(\delta^{2})$ Sec. 3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec. 3.2 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ Sec. 3.4 It depends on
the qualities of
$r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$.
[…]
Figure 1: Main results on performance gap induced by model mis-specification scenarios.
$\bullet$ When assuming exact optimization, i.e., optimization with infinite data, we study the fine-grained representation gap under different settings of reward and policy class mis-specifications in Section 3. Main results are visualized in Figure 1.
No model mis-specification: We show that the RLHF and DPO policies both achieve the performance of $\pi^{\star}$, and online DPO can further close the gap between optimization paths.
Policy model mis-specification: We show that the RLHF policy is still optimal under the model class, while the DPO policy can be sub-optimal, and online DPO cannot bridge the gap.
Reward model mis-specification: We show that the DPO policy is still optimal, while the RLHF policy can be sub-optimal due to learning based on a sub-optimal reward model.
Double model mis-specification: When policy and reward model classes are isomorphic, then they should have identical performance, while online DPO can outperform both of them. Otherwise, there is no consistent performance gap, and the comparison result depends on the qualities of (surrogate) reward models. We also give a preliminary guide for reward learning under mis-specifications.
$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the implicit representation gap incurred by statistical efficiencies in Section 4. We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$. Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\mathcal{O}(\sqrt{k\log d/n})$. This result indicates that DPO is less sample-efficient than RLHF, leading to inferior performance.
Finally, we use empirical experiments to corroborate these theoretical findings in Section 5.
2 Preliminaries
[…]
Multi-armed Bandits and Policies. A multi-armed bandit is defined by an action space $\mathcal{Y}$ and a reward function $r:\mathcal{Y}\to\mathbb{R}$. A policy $\pi\in\Delta(\mathcal{Y})$ represents a probability distribution over the action space. Given the autoregressive mechanism of LM, it is natural to consider the prompt and response together and view the combination as a whole action. In this study, our results are presented in the context of multi-armed bandits.
Model class and value function. Let =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d_{R}}\}$ denote the reward model class, and $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d_{P}}\}$ denote the policy model class, where $d_{R},d_{P}\in\mathbb{N}$. For a reward function $r$ and policy $\pi$, we define the regularized value function as:
$V_{r}^{\pi}:=\underset{y\sim\pi}{\mathbb{E}}[r(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;,$
where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$. A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, where $Z:=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$ is the partition function. The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. We define the oracle value as $V^{\star}_{\Pi}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
Bradley-Terry (BT) model. Given an implicit reward oracle $r:\mathcal{Y}\to\mathbb{R}$, Bradley and Terry (1952) assume that human preference distribution $p^{\star}:\mathcal{Y}\times\mathcal{Y}\rightarrow\Delta(\{0,1\})$ satisfies:
$p^{\star}(y_{1}\succ y_{2})=\sigma\left(r^{\star}(y_{1})-r^{\star}(y_{2})\right)\;.$
This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2})\;$ by human annotators.
Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the population MLE objective:
$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;.$
In practice, people can have a pre-collected human preference dataset $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where in the $i^{\text{th}}$ sample, $y_{w}^{(i)}\succ y_{l}^{(i)}$ means that $y_{w}^{(i)}$ is preferred over $y_{l}^{(i)}$. And the empirical loss is:
$\mathcal{L}_{\textup{RLHF}}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\;.$
Then using the reward model $r_{\textup{RLHF}}$, the policy learning stage returns $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\textrm{RLHF}}}^{\pi}$.
Direct approach of DPO. By leveraging the surrogate reward $\hat{r}_{\theta}(y):=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
We also consider an online variant of DPO, where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy. Online DPO then minimizes the modified loss:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
3 Exact Optimization: Fine-grained Representation Gap Induced by Model Mis-specification
We analyze the behavior of RLHF and DPO in the idealized setting of exact optimization, where both methods have access to infinite preference data and can optimize their respective objectives without statistical or computational error. Our focus is on the performance gap induced by the representation gap – that is, the difference between the true optimal policy and the best policy each method can produce, as determined by the expressiveness of the reward and policy model classes.
Recall that $r_{\textup{RLHF}}\in\mathcal{F}$ is the solution computed by exact optimization of reward learning, $\pi_{\textup{RLHF}}\in\Pi$ is the solution computed by exact optimization of policy learning given $r_{\textup{RLHF}}$, and $\pi_{\textup{DPO}}\in\Pi$ is the solution computed by exact optimization of DPO.
3.1 No Model Mis-specification
We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes. Most prior works (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025) focus on this setting.
Condition 1 (Strong Reward Model, Strong Policy Model).
$r^{\star}\in\mathcal{F}$, $\pi^{\star}\in\Pi$.
Both RLHF and DPO are capable of recovering the true optimal policy under ideal conditions. In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage. Proof deferred to Section C.1.
Proposition 1.
Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$.
Shi et al. (2025) propose a sampling strategy to accelerate convergence in online DPO, and Feng et al. (2025) further refine this approach, showing its connection to the RLHF objective from a gradient-based perspective. Below, we show a result which is analogous to Theorem 4.1 in (Feng et al., 2025), but from the objective perspective rather than the gradient perspective.
Definition 1 (PILAF Sampler (Shi et al., 2025; Feng et al., 2025)).
PILAF Sampler is a probabilistic mixture of two sampler pairs:
[…]
with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;$.
Remark 1.
To sample a pairwise data $(y_{1},y_{2})$ from PILAF sampler, we sample $y_{1}\sim\pi^{\mathsf{s1}},y_{2}\sim\pi^{\mathsf{s2}}$, w.p. $\alpha_{1}/(\alpha_{1}+\alpha_{2})$ from sampler ➀, and w.p. $\alpha_{2}/(\alpha_{1}+\alpha_{2})$ from sampler ➁.
Theorem 2.
Given $R_{\max},\delta\in\mathbb{R}_{+}$, s.t. $0\leq r^{\star}(y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\delta$, $\forall y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;,$
where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t. $|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z(\theta):=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$ can be viewed as an adaptive step size.
Remark 2.
This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value objective up to a second-order deviation. However, this deviation can become substantial when the reward function has a large scale or the ground-truth reward is poorly fitted. In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence. Proof deferred to Section C.2.
3.2 Policy Model Mis-specification
We now examine the setting where the ground-truth reward function is realizable ($r^{\star}\in\mathcal{F}$), but the optimal policy is non-realizable by the policy class ($\pi^{\star}\notin\Pi$). This case can be referred to Nika et al. (2024), who point out that the optimal policy could be more complicated than the optimal reward, and Swamy et al. (2025), who attribute this scenario to generation-verification gaps in fine-tuning.
Condition 2 (Strong Reward Model, Weak Policy Model).
$r^{\star}\in\mathcal{F}$, $\pi^{\star}\notin\Pi$.
In this case, RLHF has a structural advantage: it can recover the exact reward and then compute the best possible policy within $\Pi$. In contrast, DPO bypasses reward modeling and directly learns a policy from preferences, which may lead to sub-optimal behavior due to mismatches between preference-based objectives and reward-based value functions. The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization. Proof deferred to Section C.3.
Proposition 3.
Under Condition 2, $V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section C.8.
Proposition 4.
Under Condition 2, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi^{\textup{online}}_{\textup{DPO}}}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi^{\textup{online}}_{\textup{DPO}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}$ where the online sampler is PILAF sampler (Definition 1).
3.3 Reward Model Mis-specification
We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$. This setting is studied for completeness. As discussed in Swamy et al. (2024), two-stage RLHF can only lose information during reward learning, which will be highlighted under reward model mis-specification.
Condition 3 (Weak Reward Model, Strong Policy Model).
$r^{\star}\notin\mathcal{F}$, $\pi^{\star}\in\Pi$.
In this setting, RLHF is vulnerable to reward mis-specification: the learned mis-specified reward model $r_{\textup{RLHF}}$ could significantly deviate from the ground-truth reward $r^{\star}$, causing the subsequent policy optimization to yield a sub-optimal solution even though $\pi^{\star}\in\Pi$. Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error. Proof deferred to Section C.4.
Proposition 5.
Under Condition 3, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
3.4 Double Model Mis-specification
We now consider the most challenging setting, where neither the ground-truth reward function nor the optimal policy is realizable by their respective model classes.
Condition 4 (Weak Reward Model, Weak Policy Model).
$r^{\star}\notin\mathcal{F}$, $\pi^{\star}\notin\Pi$.
To enable a fine-grained comparison between RLHF and DPO under this double mis-specified regime, we introduce the surrogate reward model class induced by the policy class as $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\theta\in\mathbb{R}^{d_{P}},\hat{r}_{\theta}(y)=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)},\forall y\in\mathcal{Y}\}$. Pairwise preferences depend only on reward differences, so reward functions $r,r^{\prime}:\mathcal{Y}\to\mathbb{R}$ are equivalent if they differ by a constant. We compare the expressiveness of the original reward model class $\mathcal{F}$ and the surrogate class $\mathcal{F}_{\Pi}$, modulo constant shifts, and analyze three representative regimes characterizing their relative capacities:
Condition 5 (Isomorphism).
$r^{\star}\notin\mathcal{F}$, $\pi^{\star}\notin\Pi$. $\mathcal{F}=\mathcal{F}_{\Pi}$.
Condition 6 (Policy Model Class Is Relatively Stronger).
$r^{\star}\notin\mathcal{F}$, $\pi^{\star}\notin\Pi$. $\mathcal{F}\subset\mathcal{F}_{\Pi}$.
Condition 7 (Reward Model Class Is Relatively Stronger).
$r^{\star}\notin\mathcal{F}$, $\pi^{\star}\notin\Pi$. $\mathcal{F}\supset\mathcal{F}_{\Pi}$.
Remark 3.
Note that certain cases involve partially overlapping model classes. However, we do not consider these intermediate regimes for the sake of a principled analysis.
For Condition 5, we focus on the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies. This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage. Then motivated by Nika et al. (2024), we can split the performance of RLHF and DPO as (see detailed calculations in Section C.12):
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
In two-stage RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling. Therefore, two-stage RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO. Proofs deferred to Sections C.5 and C.9.
Proposition 6.
Under Condition 5, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Proposition 7.
Under Condition 5, there exists an environment where online DPO can produce a solution $\pi_{\textup{DPO}}^{\textup{online}}$, s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections C.6 and C.7.
Proposition 8.
Under Condition 6, there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and another environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Proposition 9.
Under Condition 7, there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and another environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Though there is no consistent performance gap between RLHF and DPO in certain settings, revisiting the framework can reveal a structural parallel: RLHF can yield the best policy given the learned reward model $r_{\textup{RLHF}}$, and the DPO policy is directly the optimal one given the surrogate reward model $\hat{r}_{\textup{DPO}}$. And online DPO serves to enhance the quality of $\hat{r}_{\textup{DPO}}$ (Xiong et al., 2024). Formally,
$\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\textup{RLHF}}}^{\pi}\;,\;\pi_{\textup{DPO}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{\hat{r}_{\textup{DPO}}}^{\pi}\;.$ (1)
This result implies a general principle: to measure the performance gap, we only need to measure the quality gap between the (surrogate) reward models.
As revealed in Appendix C of Ouyang et al. (2022) and Section 3.3 of Swamy et al. (2025), it is uncommon to deploy a reward model with greater scale than the policy model. To ensure practical relevance, we focus on the regime where $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
What key property enables a (surrogate) reward model to subsequently help learn good policies?
As an answer to this question, we can show that in the context of preference-based policy learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference: $\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}$, where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$. The detailed calculations and further discussions are deferred to Appendix B. In Section 4, we adopt a data-dependent proxy for this notion: the empirical $\ell_{2}$ distance over observed preference pairs (see Definition 2 in Section C.10, also in Zhu et al. (2023)). Using this metric, we establish a separation result that demonstrates the statistical advantage of explicitly learning reward functions.
4 Approximate Optimization: Implicit Representation Gap Induced by Statistical Efficiency in Sparse Recovery
Under more realisitc finite-sample regime, statistical efficiency can also induce performance gap. With limited preference data, we are not able to directly compute $r_{\textup{RLHF}}$, $\pi_{\textup{RLHF}}$ and $\pi_{\textup{DPO}}$, and thus obtain weaker solutions due to estimation error. This scenario can be viewed as inducing an implicit representation gap, whose effects have been widely discussed in Section 3. Motivated by Nika et al. (2024) and still assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance of two-stage RLHF as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}\leq\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{implicit reward representation gap}}\;,$
and for DPO we have (see detailed calculations in Section C.13):
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $\leq\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
We now compare the statistical efficiencies of RLHF and DPO, focusing on how their sample complexities differ and influence the implicit reward representation gap. We will further improve the analysis of Nika et al. (2024); Swamy et al. (2025), to prove that when the reward model and the policy model share the same model size, i.e. $d_{P}=d_{R}$ and even under realizability condition, two-stage RLHF is still more sample-efficient than DPO.
Difference in token-level parameterization. The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value. Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta}(y)=\beta\sum_{t=0}^{N}\theta_{t}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and $N=|y|-1$. While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al. (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
$\pi_{\theta}(y)=\prod_{t=0}^{N}\pi_{\theta}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{N}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{N}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$. Then the optimal solution for the reward model is $\theta_{t}=\theta^{\star}_{t}$, $\forall t\in\mathbb{Z}_{\geq 0}$. And the optimal solution for the policy model is given by Rafailov et al. (2024). Here we interpret the result in a slightly different way to separate $\pi_{{\mathsf{ref}}}$ and the $q$ function (see detailed calculations in Section C.13):
$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$ (2)
where the $q$ function is determined in a recursive way:
$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
Benefiting from the token-level $q$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025). However, simultaneously, learning the $q$ function sacrifices statistical efficiency due to the need to model the complicated structure. Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
Dual-token Sparse Prediction (DTSP).  Under log-linear policy class, the policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
We let $\theta_{r}^{\star}$ denote the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ the optimal solution for surrogate reward learning. Note that for the second token, $\theta_{r}^{\star}$ and $\theta_{p}^{\star}$ share the same optimal solution:
$(\theta_{r,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{1}\;,\;(\theta_{p,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{2}\;,$
where $C_{1},C_{2}\in\mathbb{R}$ are offsets. And for the first token $y$, there is a distinction:
$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
where $\textbf{r}_{\textup{sparse}}$ gets entangled with $\pi_{{\mathsf{ref}}}$ in $\theta_{p,0}^{\star}$. Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2. And even without model mis-specification, we can show that an implicit representation gap can still arise due to statistical inefficiencies.
Theorem 10 (Efficient Sparse Recovery).
For DTSP task, and under mild assumptions, there exists an environment, s.t. by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(k\log d/n)$ using an $\ell_{0}$-constrained estimator, and can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator.
Theorem 11 (Separation).
For DTSP task, and under mild assumptions, there exists an environment, s.t. by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using an $\ell_{1}$-regularized estimator, while the estimation error of the policy model is lower bounded by $\Omega(d/n)$.
Remark 4.
Benefiting from fixing optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to one-token prediction, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense. Pure reward learning can thus reduce the estimation error to $\mathcal{O}(1)$ using $\mathcal{O}\left(k\log d\right)$ samples instead of $\mathcal{O}(d)$ samples. Detailed proofs deferred to Section C.10.
Takeaways. This section shows that the policy model is more prone to model mis-specification than the reward model, under a realistic setting, and thus Conditions 1 and 2 are more prevalent than Condition 3. And even without model mis-specification, the estimation error can also induce an implicit representation gap. From the perspective of sparse recovery, we can see that the surrogate reward model could suffer from severe statistical inefficiency, while the reward model can recover the reward structure very efficiently, even with the same model scale. Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model. This further explains why two-stage RLHF is empirically observed to outperform DPO (Ivison et al., 2024; Xu et al., 2024b).
5 Experimental Verifications
Experiment setup. We now verify our analysis in practical settings. We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023). We first finetune a GPT-2-large-774M model (Radford et al., 2019) on $5$k samples of PKU-SafeRLHF-QA, and obtain the SFT model. We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle. All experiments are repeated for $3$ seeds. Please see Appendix D for more details.
Implementation details. For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset. For approximate optimization, we instead compute the empirical BT loss. We adopt a pairwise variant of policy gradient instead of PPO to improve training stability: $\mathcal{L}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1}),\hat{r}_{\theta}(y_{2}))\right]^{2}$. During deployment, the ground-truth reward will be scaled by a coefficient $r_{\textup{margin}}$. Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
Verifications of Section 3. We train online DPO and RLHF on PKU-SafeRLHF-Prompt, following the practice of Dong et al. (2024); Shi et al. (2025). For the strong reward condition, we directly adopt the GPT2-large-harmless model as a perfectly-learned reward model. For the weak reward condition, we train the SFT model on PKU-SafeRLHF-safer by replacing the projection matrix with a linear head, freezing all layers except the linear head and the last block. For the strong policy condition, we fully train the SFT model, while for the weak policy condition, we freeze the first half of the blocks of the SFT model. Results are shown in Figures 2 and 3.
Verifications of Section 4. We train offline DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al. (2024), and we train on two types of preference: “better” and “safer”, and down-sampled the corresponding training datasets to $1$k-$9$k samples. For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head. The models are trained under the same setting, and achieve at least $85\%$ training accuracy. Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
Figure 2: Experimental Results for Condition 1. We experiment on different reward scales $0.4,1,4$, and observe that the advantage of RL over online DPO is larger for larger reward scales.
Figure 3: Experimental Results for Conditions 2, 3 and 4. The first two figures align with our claims. And the gap in the last figure can be attributed to the mis-specified reward model being too weak.
Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types, and pure reward learning is shown to be more data-efficient than surrogate reward learning.
6 Related Works
Performance gap between RLHF and DPO. Recently, there have been works investigating the performance gap between RLHF and DPO policies. Xu et al. (2024b) found that DPO might find biased solutions that exploit out-of-distribution responses, and iterative DPO might be a better approach; meanwhile, PPO with advantage normalization, large batch-size, and exponential moving update of the reference model can consistently outperform DPO on benchmarks.
Swamy et al. (2025) first showed that when the reward class and policy class are isomorphic, RLHF and DPO output policies with equal performances. Then, they proposed a hypothesis that when the ground-truth reward is simpler than the soft optimal policies, and the reward class reduces the sample complexity to learn such a reward, then reward modeling essentially reduces the policy search space. This hypothesis is supported by their experiments. In our work, we comprehensively extend upon their first class isomorphic result by studying model mis-specification (Section 3), and we construct concrete examples to further support the existence of the “simpler ground-truth reward” and “reduced sample complexity” (Section 4).
Nika et al. (2024) provided sub-optimality upper bounds for RLHF and DPO when assuming linear reward class and log-linear policy class, with the un-regularized value as performance metric. Three cases were studied: ➀ realizable ground-truth reward and exact optimization, ➁ realizable ground-truth reward but approximate optimization, as well as ➂ non-realizable reward and exact optimization. Let $n$ be the size of the fixed dataset and $d$ be the feature dimension. For case ➀, both algorithms have a policy bias due to the un-regularized metric, while RLHF has an additional $\Theta(\sqrt{d/n})$ statistical error and that for DPO is $\Theta(d/(\beta n))$. For case ➁, RLHF and DPO both obey a linear convergence to statistical errors and policy biases when using projected gradient descent. For case ➂, aside from statistical errors and policy biases, RLHF has an extra approximation error between the ground-truth reward and best achievable reward, while DPO has an extra bias between the optimal regularized policy and the ideal optimal regularized policy.
Now we focus on comparing with the most relevant prior study, Nika et al. (2024). First, unlike Nika et al. (2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: ➀ it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; ➁ it can help circumvent the unavoidable upper bound of policy bias in the unregularized version. Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes. Third, we improve the statistical analysis of Nika et al. (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\mathcal{O}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
A more general list of related works is deferred to Appendix A.
7 Conclusion
This paper provides a fine-grained analysis of the performance gap between two-stage and direct approaches to preference-based policy learning. We theoretically demonstrate a dichotomy of RLHF and DPO under different mis-specification scenarios, and further reveal a implicit representation gap induced by statistical efficiency. Our claims are supported by empirical experiments on LMs.
It is also important to acknowledge our limitations. ➀ While we identify a limitation of training reward models based on BT model, we do not provide a theoretically grounded and practically effective alternative. ➁ Due to computational constraints, our experiments are limited to small-scale models. We hope our insights can motivate the community to further investigate these directions.
Acknowledgement
SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P. Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship. RZ and MF acknowledge the support of NSF TRIPODS II DMS-2023166. The work of MF was supported in part by awards NSF CCF 2212261 and NSF CCF 2312775. RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
Appendix
[…]
Appendix A Supplementary Related Works
Reinforcement learning from human feedback (RLHF). Seminal contributions that showcased RLHF’s applicability to LLMs include foundational work by Christiano et al. [2017], and subsequent research focusing on tasks such as summarization [Stiennon et al., 2020], instruction following [Ouyang et al., 2022], question answering using web-retrieved information [Nakano et al., 2021], and broader AI alignment objectives [Bai et al., 2022]. Theoretical studies of RLHF include pessimism in policy learning [Zhu et al., 2023], overoptimization [Zhu et al., 2024, Liu et al., 2024c], online RLHF [Xiong et al., 2024, Song et al., 2024], robustness [Mandal et al., 2025], and reward models [Wang et al., 2024, Razin et al., 2025b, Huang et al., 2025, Yao et al., 2025].
Direct preference optimization (DPO). There is a rich literature studying offline [Rafailov et al., 2024, Feng et al., 2024], iterative [Dong et al., 2024, Liu et al., 2024a], and online [Guo et al., 2024, Tajwar et al., 2024, Ding et al., 2024, Shi et al., 2025, Chen et al., 2025, Feng et al., 2025] DPO. There are other DPO-style algorithms to directly optimize the policy model from preference signals, such as $\Psi$-PO [Azar et al., 2023], RSO [Liu et al., 2024b], RS-DPO [Khaki et al., 2024], CPO [Xu et al., 2024a], SimPO [Meng et al., 2024], XPO [Xie et al., 2024], VPO [Cen et al., 2024], and OAIF [Guo et al., 2024].
Appendix B Bonus: How Can We Better Model Reward From Preference Signals?
As motivated by Equation 1, a reward model $r_{\phi}$ can be mapped to a policy via:
$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\phi}}^{\pi}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;\underset{y\sim\pi}{\mathbb{E}}[r_{\phi}(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;.$
If $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, this solution further admits the closed form $\pi_{\theta^{\star}(r_{\phi})}(y)=\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)/Z(\phi)$, where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)$ is the partition function. If the goal is to output a policy that performs well under the ground-truth reward $r^{\star}$, then reward learning should aim to find a model $r_{\phi}$ such that the resulting policy $\pi_{\theta^{\star}(r_{\phi})}$ maximizes the underlying “real” objective:
$r_{\phi^{\star}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmin}}\;\underbrace{-\beta\log Z(\phi)-\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:\mathcal{L}_{\textup{new}}(\phi)}\;.$
Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section C.12):
$\nabla_{\phi}\mathcal{L}_{\textup{new}}(\phi)=-\frac{1}{2}\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;,$ (3)
which corresponds to the gradient of an $\ell_{2}$ distance of pairwise difference:
$\mathcal{L}_{\textup{new}}(\phi){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{4}\,\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}{\mathbb{E}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;.$ (4)
Comparison with MLE. The reward model $r_{\phi}$ are typically learned via MLE from preference data, which does not consider the fact that the learned reward will ultimately be used to induce a policy. Now revisit the MLE objective:
$\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+\sigma(r^{\star}(y^{\prime})-r^{\star}(y))\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;,$
whose gradient is:
$\nabla_{\phi}\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;.$ (5)
Comparing Equation 3 with Equation 5, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 5 to further align with Equation 3. And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$. And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\propto\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$. If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it is a process reward model. However, it is worth noting that the model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2. Therefore, when faced with representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
To alleviate this issue, the remedy could be to learn the preference with alternative modeling approaches to circumvent the BT model setting, which has already be revealed in the successes of Sun et al. [2025], Calandriello et al. [2024]. For example, we can look into the training objective of online IPO [Calandriello et al., 2024, Zhou et al., 2025] (see detailed calculations in Section C.12):
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y_{1},y_{2})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))-\frac{p^{\star}(y_{1}\succ y_{2})-p^{\star}(y_{2}\succ y_{1})}{2}\right]^{2}\;,$
where $\rho(\theta)$ is an online sampling distribution, and it thus can optimize a $\ell_{2}$ distance in an online way. The classification model deployed in Sun et al. [2025] is also promising. We leave this interesting direction for future exploration, and now come back to two-stage RLHF and DPO.
Appendix C Omitted Proofs
C.1 Proof of Proposition 1
Since $r^{\star}\in\mathcal{F}$, RLHF exactly recovers $r^{\star}$ during reward learning. The policy optimization stage then solves $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi}$, so by definition, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$.
On the other hand, DPO is trained on preferences induced by $r^{\star}$. When $\pi^{\star}\in\Pi$, the preference structure is realizable, and the DPO loss is minimized by $\pi^{\star}$. Hence, $\pi_{\textup{DPO}}=\pi^{\star}$, which achieves the maximum of $V_{r^{\star}}^{\pi}$ over $\Pi$.
C.2 Proof of Theorem 2
By Taylor’s expansion, we have that:
$\nabla_{\theta}\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ $=-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]$
$\quad-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime\prime}(\xi_{y,y^{\prime}})\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]^{2}\;,$
where $\xi_{y,y^{\prime}}$ is an intermediate value between $r^{\star}(y)-r^{\star}(y^{\prime})$ and $\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})$.
Therefore, if we have:
  • $0\leq r(y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$;
  • $|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\delta$, $\forall y,y^{\prime}\in\mathcal{Y}$;
  • $\pi^{\mathsf{s}}(y,y^{\prime})\propto\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, i.e., $\pi^{\mathsf{s}}$ is PILAF sampler,
then the formula can be rewritten as:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{2\mathsf{sg}\left(Z(\theta)\right)}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left(1+\epsilon_{y,y^{\prime}}\right)\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;,$
where
$|\epsilon_{y,y^{\prime}}|$ $=\left|\frac{\sigma^{\prime\prime}(\xi_{y,y^{\prime}})}{\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))}\right|\cdot|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|$
$\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime}(R_{\max}+\delta)}\;,$
and
$Z(\theta):=\sum_{y,y^{\prime}\in\mathcal{Y}}\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;.$
Note that:
$\nabla_{\theta}V_{r^{\star}}^{\pi_{\theta}}$ $=\nabla_{\theta}\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(y)-\hat{r}_{\theta}(y)]$
$=\underset{y\sim\pi_{\theta}}{\mathbb{E}}\nabla_{\theta}\log\pi_{\theta}(y)[r^{\star}(y)-\hat{r}_{\theta}(y)]$ (policy gradient theorem)
$=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\nabla_{\theta}\log\pi_{\theta}(y)[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})]$ (policy gradient theorem)
$=\frac{1}{2}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right][(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})]\;,$ (symmetry)
thus
$V_{r^{\star}}^{\pi_{\theta}}$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}-\frac{1}{4\beta}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;.$
Therefore we have:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;.$
C.3 Proof of Proposition 3
Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$. DPO, instead, minimizes a proxy loss defined over pairwise preferences. Since $\pi_{\textup{DPO}}\in\Pi$, we have $V_{r^{\star}}^{\pi_{\textup{DPO}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}$, which proves the first claim.
For the strict gap, we consider a multi-armed bandit setting with the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
$r=r^{\star}(a_{1})=r^{\star}(a_{2})\geq r^{\star}(a_{3})=0\;.$
Assume the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{d}$ satisfies:
$\psi(a_{1})\neq\psi(a_{2})\;,\;\psi(a_{3})=\tfrac{1}{2}\psi(a_{1})+\tfrac{1}{2}\psi(a_{2})\;.$
Define the log-linear policy class $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d}\}$ by $\pi_{\theta}(a)\propto\pi_{{\mathsf{ref}}}(a)\exp(\theta^{\top}\psi(a))$, where $\pi_{{\mathsf{ref}}}=\mathrm{Unif}(\mathcal{Y})$. Since $r^{\star}$ is realizable, RLHF exactly recovers it and solves:
$\pi_{\textup{RLHF}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi_{\theta}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\sum_{a\in\mathcal{Y}}\pi_{\theta}(a)r^{\star}(a)-\beta\,\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\;.$
For a fixed $r>0$, as the regularization parameter $\beta\to 0$, the optimal policy under RLHF places vanishing probability on $a_{3}$: $\pi_{\textup{RLHF}}(a_{3})\to 0$. In contrast, as $\beta\to\infty$, the regularization dominates and the optimal policy converges to the uniform reference policy: $\pi_{\textup{RLHF}}\to\pi_{{\mathsf{ref}}}$.
Now consider the DPO objective, which relies on pairwise preference probabilities and directly optimizes over the policy class:
$\mathcal{L}_{\textup{DPO}}(\theta)$ $=-\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma\left(\beta\,\theta^{\top}(\psi(a)-\psi(a^{\prime})\right)\right]$
$=-\tfrac{1}{2}\log\sigma(\beta\Delta^{\top}\theta)-\tfrac{1}{2}\log\sigma(-\beta\Delta^{\top}\theta)-\log\sigma(\tfrac{1}{2}\beta\Delta^{\top}\theta)-\log\sigma(-\tfrac{1}{2}\beta\Delta^{\top}\theta)\;,$
where $\Delta:=\psi(a_{1})-\psi(a_{2})$. This expression is always minimized when $\Delta^{\top}\theta=0$, which corresponds to a uniform distribution.
Thus, unlike RLHF, the DPO solution remains fixed at uniform distribution, independent of the reward magnitude $r$ and the regularization parameter $\beta$, and fails to suppress the sub-optimal action $a_{3}$ even when $\beta$ is sufficiently small.
C.4 Proof of Proposition 5
Since $r^{\star}\notin\mathcal{F}$, RLHF recovers an approximation $\hat{r}\in\mathcal{F}$ via reward learning. It then computes a policy $\pi_{\textup{RLHF}}$ that maximizes $V_{\hat{r}}^{\pi}$ over $\Pi$. In general, this policy is sub-optimal under $r^{\star}$ (see Proposition 3), and thus $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}$.
DPO directly optimizes a preference-based loss over $\Pi$. Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V_{\Pi}^{\star}$.
For the strict gap, consider a multi-armed bandit setting analogous to Section C.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
$r=r^{\star}(a_{1})=r^{\star}(a_{2})\geq r^{\star}(a_{3})=0\;.$
Assume the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{d}$ satisfies:
$\psi(a_{1})\neq\psi(a_{2})\;,\;\psi(a_{3})=\tfrac{1}{2}\psi(a_{1})+\tfrac{1}{2}\psi(a_{2})\;.$
The key difference from the earlier construction lies in the choice of model classes. We define: the linear reward class =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d}\}$ by $r_{\phi}(a):=\phi^{\top}\psi(a)$, and the policy class $\Pi=\Delta(\mathcal{Y})$ with reference policy $\pi_{{\mathsf{ref}}}=\textrm{Unif}(\mathcal{Y})$. This setup satisfies Condition 3 because $r^{\star}\notin\mathcal{F}$: for any $\phi$, the constraint on $\psi$ implies $r_{\phi}(a_{3})=\tfrac{1}{2}(r_{\phi}(a_{1})+r_{\phi}(a_{2}))$ so $r_{\phi}(a_{3})=r$ whenever $r_{\phi}(a_{1})=r_{\phi}(a_{2})=r$, contradicting the ground-truth reward $r^{\star}(a_{3})=0$.
In RLHF, the reward model is learned by solving the population MLE objective:
$r_{\textup{RLHF}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma(\beta\phi^{\top}(\psi(a)-\psi(a^{\prime})))\right]$
$=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;-\tfrac{1}{2}\log\sigma(\beta\Delta^{\top}\phi)-\tfrac{1}{2}\log\sigma(-\beta\Delta^{\top}\phi)-\log\sigma(\tfrac{1}{2}\beta\Delta^{\top}\phi)-\log\sigma(-\tfrac{1}{2}\beta\Delta^{\top}\phi)\;,$
where $\Delta:=\psi(a_{1})-\psi(a_{2})$. This expression is minimized at $\Delta^{\top}\phi=0$, which implies $r_{\phi}(a_{1})=r_{\phi}(a_{2})$ and $r_{\phi}(a_{3})=r_{\phi}(a_{1})$, i.e., the learned reward is constant: $r_{\textup{RLHF}}(a)=C$ for all $a\in\mathcal{Y}$.
The policy learning stage then solves:
$\pi_{\textup{RLHF}}=\underset{\pi\in\Delta(\mathcal{Y})}{\operatorname{argmax}}\;\underset{a\sim\pi}{\mathbb{E}}[C]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right))\;,$
whose solution is simply $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
In contrast, DPO directly optimizes the policy using preference comparisons. Since $\Pi=\Delta(\mathcal{Y})$ and the preferences are consistent with the ground-truth reward $r^{\star}$, DPO can recover the optimal policy $\pi^{\star}\propto\exp(r^{\star}/\beta)$, which is not uniform. Therefore, DPO achieves the optimal regularized value $V^{\star}_{\Pi}=V_{r^{\star}}^{\pi^{\star}}$, while RLHF only returns the uniform policy. This yields a strict gap:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}\;.$
C.5 Proof of Proposition 6
By definition, the reward learned by RLHF and the surrogate reward learned by DPO are obtained by solving the following population objectives:
$r_{\textup{RLHF}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;,$
$\hat{r}_{\textup{DPO}}$ $=\underset{\hat{r}_{\theta}\in\mathcal{F}_{\Pi}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y))\right]\;,$
Under Condition 5, we have $\mathcal{F}=\mathcal{F}_{\Pi}$ , so both objectives are optimized over the same function class. Hence, it follows that: $r_{\textup{RLHF}}=\hat{r}_{\textup{DPO}}$.
Recalling from Equation 1:
$\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\textup{RLHF}}}^{\pi}\;,\;\pi_{\textup{DPO}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{\hat{r}_{\textup{DPO}}}^{\pi}\;.$
and substituting $r_{\textup{RLHF}}=\hat{r}_{\textup{DPO}}$, we can conclude that
$\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}\quad\text{and hence}\quad V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
C.6 Proof of Proposition 8
Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. We first construct an environment satisfying Condition 6 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider the same setup as in Section C.4, but define the policy class as $\Pi=\Delta(\mathcal{Y})\setminus\{\pi^{\star}\}$, where $\pi^{\star}$ is the optimal policy under $r^{\star}$. This ensures that $\pi^{\star}\notin\Pi$, while $\mathcal{F}\subset\mathcal{F}_{\Pi}$, satisfying Condition 6.
As shown in Section C.4, RLHF learns a constant reward model and returns the uniform policy $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$. In contrast, DPO directly optimizes policy parameters from preference data and can converge to a policy arbitrarily close to $\pi^{\star}$, which lies on the boundary of $\Pi$. This yields a strict sub-optimality gap:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
Construction 2: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Next, we construct an environment satisfying Condition 6 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward:
$r^{\star}(a_{1})=r^{\star}(a_{2})=1\;,\;r^{\star}(a_{3})=0\;.$
Let the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{2}$ be:
$\psi(a_{1})=\begin{bmatrix}1\\ 0\end{bmatrix},\;\psi(a_{2})=\begin{bmatrix}0\\ 1\end{bmatrix},\;\psi(a_{3})=\begin{bmatrix}1/2\\ 1/2\end{bmatrix}.$
Define the log-linear policy class $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{2}\}$ with
$\pi_{\theta}(a)\propto\pi_{{\mathsf{ref}}}(a)\exp(\theta^{\top}\psi(a))\;,\;\pi_{{\mathsf{ref}}}=\mathrm{Unif}(\mathcal{Y}).$
The corresponding surrogate reward class is $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a),\;\theta\in\mathbb{R}^{2}\}$. We now define a strictly smaller reward model class =$\mathcal{F}=\{\hat{r}_{\theta_{R}}\}$ where
$\theta_{R}=\begin{bmatrix}1\\ -1\end{bmatrix}.$
We set the regularization parameter to $\beta=0.1$. Then, $\mathcal{F}\subset\mathcal{F}_{\Pi}$ and Condition 6 holds.
Under this setup, RLHF learns the fixed reward $\hat{r}_{\theta_{R}}$ and optimizes:
$\pi_{\textup{RLHF}}=\pi_{\theta_{R}}\;,\;\text{where }\pi_{\theta_{R}}(a)\propto\exp(\theta_{R}^{\top}\psi(a))\;.$
Concretely:
$\pi_{\theta_{R}}(a_{1})=\frac{\exp(1)}{Z}\;,\;\pi_{\theta_{R}}(a_{2})=\frac{\exp(-1)}{Z}\;,\;\pi_{\theta_{R}}(a_{3})=\frac{1}{Z}\;,\;Z=\exp(1)+\exp(-1)+1\;.$
The value of this policy under $r^{\star}$ is:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}=\pi_{\theta_{R}}(a_{1})+\pi_{\theta_{R}}(a_{2})-\beta\,\mathrm{KL}(\pi_{\theta_{R}}\|\pi_{{\mathsf{ref}}})\approx 0.729\;.$
In contrast, DPO learns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, as shown in Section C.3. Its value is:
$V_{r^{\star}}^{\pi_{\textup{DPO}}}=\frac{2}{3}\;.$
This results in a strict sub-optimality gap in the opposite direction:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
C.7 Proof of Proposition 9
Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$. We construct an environment satisfying Condition 7 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward function:
$r^{\star}(a_{1})=r^{\star}(a_{2})=1\;,\;r^{\star}(a_{3})=0\;.$
Let the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{2}$ be:
$\psi(a_{1})=\begin{bmatrix}1\\ 0\end{bmatrix},\;\psi(a_{2})=\begin{bmatrix}0\\ 1\end{bmatrix},\;\psi(a_{3})=\begin{bmatrix}1/2\\ 1/2\end{bmatrix}\;.$
Define the log-linear policy class $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{2}\}$ with:
$\pi_{\theta}(a)\propto\pi_{{\mathsf{ref}}}(a)\exp(\theta^{\top}\psi(a))\;,\;\pi_{{\mathsf{ref}}}=\mathrm{Unif}(\mathcal{Y})\;.$
The corresponding surrogate reward class is $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a),\;\theta\in\mathbb{R}^{2}\}$. Now define a strictly larger reward model class:
$\mathcal{F}=\mathcal{F}_{\Pi}\cup\{\bar{r}\}\;,\;\text{where }\bar{r}(a_{1})=\bar{r}(a_{2})=2,\;\bar{r}(a_{3})=0\;.$
Then $\mathcal{F}_{\Pi}\subset\mathcal{F}$, and thus Condition 7 holds.
From Section C.3, we know that DPO learns a constant reward model under this feature structure and returns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
RLHF, on the other hand, optimizes the MLE objective over the larger class $\mathcal{F}$ and selects $\bar{r}$, which achieves a higher likelihood than any function in $\mathcal{F}_{\Pi}$. Then, the learned policy is:
$\pi_{\textup{RLHF}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;V_{\bar{r}}^{\pi_{\theta}}\;.$
As $\beta\to 0$, the optimal policy $\pi_{\textup{RLHF}}$ places vanishing mass on $a_{3}$, since $\bar{r}(a_{3})=0$ while $\bar{r}(a_{1})=\bar{r}(a_{2})=2$. Hence, $\pi_{\textup{RLHF}}(a_{3})\to 0$.
This leads to a strictly better policy under $r^{\star}$ than the uniform policy returned by DPO. Thus:
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
Construction 2: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. We now construct an environment satisfying Condition 7 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward function:
$r^{\star}(a_{1})=r^{\star}(a_{2})=1\;,\;r^{\star}(a_{3})=0\;.$
Let the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{2}$ be:
$\psi(a_{1})=\begin{bmatrix}1\\ 0\end{bmatrix},\;\psi(a_{2})=\begin{bmatrix}0\\ 1\end{bmatrix},\;\psi(a_{3})=\begin{bmatrix}1/2\\ 1/2\end{bmatrix}.$
We define a constrained log-linear policy class:
$\Pi=\left\{\pi_{\theta}:\theta\in\mathbb{R}^{2},\;\theta^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}\geq 20\right\}\;,\;\pi_{\theta}(a)\propto\pi_{{\mathsf{ref}}}(a)\exp(\theta^{\top}\psi(a))\;,$
where $\pi_{{\mathsf{ref}}}=\mathrm{Unif}(\mathcal{Y})$. The corresponding surrogate reward class is:
$\mathcal{F}_{\Pi}=\left\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a)\;,\;\theta^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}\geq 20\right\}\;.$
Now define a strictly larger reward model class:
$\mathcal{F}=\mathcal{F}_{\Pi}\cup\{\bar{r}\}\;,\;\text{where }\bar{r}(a_{1})=\bar{r}(a_{2})=2\;,\;\bar{r}(a_{3})=0\;.$
We set the regularization parameter to $\beta=0.1$. Since $\bar{r}\notin\mathcal{F}_{\Pi}$, we have $\mathcal{F}_{\Pi}\subset\mathcal{F}$, and thus Condition 7 holds.
Under this setup, RLHF first learns the reward model by optimizing the MLE objective over the larger class $\mathcal{F}$ and selects $\bar{r}$, which achieves strictly higher likelihood than any element in $\mathcal{F}_{\Pi}$. In the policy learning stage, RLHF computes the policy $\pi_{\textup{RLHF}}=\pi_{\theta_{\textup{RLHF}}}$ by solving:
$\pi_{\theta_{\textup{RLHF}}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;V_{\bar{r}}^{\pi_{\theta}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\left\{2\big{(}\pi_{\theta}(a_{1})+\pi_{\theta}(a_{2})\big{)}-\beta\,\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\right\}.$
In contrast, DPO directly optimizes the reward via MLE:
$\hat{r}_{\textup{DPO}}=\underset{\hat{r}_{\theta}\in\mathcal{F}_{\Pi}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y))\right]\;,$
whose optimal solution corresponds to $\theta$ satisfying $\theta^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}=20$. Therefore, the learned policy is $\pi_{\textup{DPO}}=\pi_{\theta_{\textup{DPO}}}$ with $\theta_{\textup{DPO}}^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}=20$.
To compare the values $V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ and $V_{r^{\star}}^{\pi_{\textup{DPO}}}$, we rewrite the value function for any $\pi_{\theta}$ as:
$V_{r^{\star}}^{\pi_{\theta}}$ $=\pi_{\theta}(a_{1})+\pi_{\theta}(a_{2})-\beta\,\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)$
[…]
where $x:=\theta^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}$ and $Z(x):=e^{x/2}+e^{-x/2}+1$.
It can be verified that $V_{r^{\star}}^{\pi_{\theta}}$ is strictly decreasing in $x$ for $x\geq 20$. Since RLHF learns $x_{\textup{RLHF}}\approx 40$ and DPO learns $x_{\textup{DPO}}=20$, we conclude that
$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}\;,$
demonstrating that a more expressive reward model class may lead RLHF to overfitting in the presence of a constrained policy class, resulting in inferior performance compared to DPO.
C.8 Numerical Proof of Proposition 4
By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$. Now we construct an environment under Condition 2, such that online DPO cannot outperform DPO, even with PILAF sampler. Consider a multi-armed bandit with action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$ and ground-truth reward:
$r^{\star}(a_{1})=12\;,\;r^{\star}(a_{2})=12\;,\;r^{\star}(a_{3})=0\;.$
Let the linear feature mapping $\psi:\mathcal{Y}\to\mathbb{R}^{d}$ satisfies:
$\psi(a_{1})\neq\psi(a_{2})\;,\;\psi(a_{3})=\frac{1}{2}\psi(a_{1})+\frac{1}{2}\phi(a_{2})\;.$
Taking $\beta=1$, let $x(\theta)$ denote $\log\frac{\pi_{\theta}(a_{1})}{\pi_{{\mathsf{ref}}}(a_{1})}-\log\frac{\pi_{\theta}(a_{2})}{\pi_{{\mathsf{ref}}}(a_{2})}$. Define the bounded log-linear policy class $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d}\;,\;|x(\theta)|\leq 4\}$ with
$\pi_{\theta}(a)\;\propto\;\pi_{{\mathsf{ref}}}(a)\exp(\theta^{\top}\psi(a))\;,\;\;\pi_{{\mathsf{ref}}}=\mathrm{Unif}(\mathcal{Y})\;.$
Note that we can use $x(\theta)$ to represent the whole distribution thanks to the feature mapping. Now we numerically compute the gradients of the loss functions of RL, DPO, and online DPO with PILAF sampler, in the interval $x(\theta)\in[-4,4]$. And the curves along with respective solutions are shown in the left panel of Figure 5, where the gradient values are rescaled for clarity of presentation. We find that both DPO and online DPO will converge to the same sub-optimal solution, while RL can obtain an optimal solution.
C.9 Numerical Proof of Proposition 7
By Proposition 6, we have $\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}$. Now we only need to construct an environment under Condition 5, such that online DPO can outperform offline DPO. We can borrow the whole setting in Section C.8, while resetting the ground-truth reward as:
$r^{\star}(a_{1})=24\;,\;r^{\star}(a_{2})=12\;,\;r^{\star}(a_{3})=0\;.$
Now we numerically compute the gradients of the loss functions of DPO and online DPO with a pure online sampler, in the interval $x(\theta)\in[-4,4]$. And the curves along with respective solutions are shown in the right panel of Figure 5, where the gradient values are rescaled for clarity of presentation. We find that online DPO can help obtain better solution than DPO, which indicates that under Condition 5, online DPO can produce a solution $\pi_{\textup{DPO}}^{\textup{online}}$, such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
Figure 5: Numerically Computed Curves of Gradient Functions and Value Functions.
C.10 Proofs of Theorems 10 and 11
Recall that to train a (surrogate) reward model, people first collect a dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$. Following BT model, $y_{1}$ is preferred over $y_{2}$, (i.e. $y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p. $\sigma(r^{\star}(y_{1})-r^{\star}(y_{2})$, where $r^{\star}(y)=(\theta^{\star})^{\top}\psi(y)$, $\theta^{\star}\in\mathbb{R}_{+}$ is the ground-truth reward vector, $\psi(y)$ is the feature vector satisfying $\|\psi(y)\|_{2}\leq L$, and $L\in\mathbb{R}_{+}$. The MLE estimator is defined as:
$\hat{\theta}_{\textup{MLE}}=\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\theta^{\top}(\psi(y_{w})-\psi(y_{l})))\;,$ (6)
where $\Theta_{B}=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $B\in\mathbb{R}_{+}$. And we assume $\theta^{\star}\in\Theta_{B}$. The data-dependent performance measure is the data-induced semi-norm [Zhu et al., 2023], which is defined as:
Definition 2 (Data-induced semi-norm).
The empirical error of an estimate $\hat{\theta}$ is defined as:
$\|\hat{\theta}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}:=\frac{1}{N}\sum_{i=1}^{N}\left[(r_{\hat{\theta}}(y_{w}^{(i)})-r_{\hat{\theta}}(y_{l}^{(i)}))-(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))\right]^{2}\;,$
where $\Sigma_{\mathcal{D}}$ is the Gram matrix:
$\Sigma_{\mathcal{D}}:=\frac{1}{N}\sum_{i=1}^{N}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}\;.$
Remark 5.
Note that $\hat{\theta}_{1}$ has been fixed as $e_{1}$ in our context, we thus only need to measure $\hat{\theta}_{0}$.
Before proceeding, we first prepare some ingredients. It is worth noting that those lemmas below only work for the one-token scenario, and we will adopt them in the dual-token prediction task later.
Lemma 1 (Theorem 1.a of Shah et al. [2015]).
For a sample size $n\geq c_{1}\text{tr($\Sigma_{\mathcal{D}}$)}$, any estimator $\hat{\theta}$ based on $n$ samples has a lower bound as:
$\underset{\theta^{\star}\in\Theta_{B}}{\sup}\mathbb{E}\left[\|\hat{\theta}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}\right]=\Omega\left(\frac{d}{n}\right)\;.$
Remark 6.
Here $c_{1}$ is a numerical constant. This lemma is to establish a information-theoretical lower bound of one-token reward learning. Therefore, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ defined in Equation 6 is lower bounded.
Lemma 2 (Lemma 3.1 of Zhu et al. [2023]).
With probability at least $1-\delta$, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ has an upper bound:
$\|\hat{\theta}_{\textup{MLE}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;.$
Definition 3 ($\ell_{0}$-constrained estimator).
$\hat{\theta}_{\ell_{0}}^{k}\in\underset{\theta\in\Theta_{B},\|\theta\|_{0}\leq k}{\operatorname{argmin}}\;\mathcal{L}_{\textup{MLE}}(\theta)\;.$
Lemma 3 (Theorem 3.2 of Yao et al. [2025]).
Consider $\|\theta^{\star}\|_{0}=k$ and $k\leq d/2$, then with probability at least $1-\delta$, the $\ell_{0}$-constrained estimator $\hat{\theta}_{\ell_{0}}^{k}$ has an upper bound:
$\|\hat{\theta}_{\ell_{0}}^{k}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
Definition 4 ($\ell_{1}$-regularized estimator).
$\hat{\theta}_{\ell_{1}}\in\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}\;\mathcal{L}_{\textup{MLE}}(\theta)+\gamma\|\theta\|_{1}\;.$
Lemma 4 (Theorem 3.3 of Yao et al. [2025]).
Consider $\|\theta^{\star}\|_{0}=k$, then with probability at least $1-\delta$, the $\ell_{1}$-regularized estimator $\hat{\theta}_{\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$ has an upper bound:
$\|\hat{\theta}_{\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\sigma)}{n}}\right)\;.$
Definition 5 (Relative $\ell_{1}$-regularized estimator).
Given $\tau\in\Theta_{B}$, the relative $\ell_{1}$-regularized estimator is defined as:
$\hat{\theta}_{\textup{rel}\ell_{1}}\in\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}\;\mathcal{L}_{\textup{MLE}}(\theta)+\gamma\|\theta-\tau\|_{1}\;.$
Lemma 5 (Generalized version of Lemma 4).
Consider $\tau\in\Theta_{B}$, $\|\theta^{\star}-\tau\|_{0}=k$, then with probability at least $1-\delta$, the relative $\ell_{1}$-regularized estimator $\hat{\theta}_{\textup{rel}\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\frac{\log 2d+\log(1/\delta)}{n}\right)$ has an upper bound:
$\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
Proof of this lemma is deferred to Section C.11.
Proof of Theorem 10. We first prepare the settings and assumptions. For the feature mapping, we set $\psi(y,\omega)=\psi(\omega)$. For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$. And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank, and $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$. Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})=\left(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector and $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
Recall that:
$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})+C_{4}\;,$
we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
Then observe the empirical MLE objective of pure reward learning:
$\hat{\theta}_{r}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{r,0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
and the empirical MLE objective of surrogate reward learning:
$\hat{\theta}_{p}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
We can thus directly apply Lemma 2 to obtain that w.p. $1-\delta$, we have:
$\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;,$
and
$\|\hat{\theta}_{r,0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;.$
Note that $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$, we can have an $\ell_{0}$-constrained estimator for the reward model:
$\hat{\theta}_{r,\ell_{0}}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k,\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{r,\ell_{0},0}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
then Lemma 3 implies that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\ell_{0},0}^{k}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
And we also have an $\ell_{1}$-regularized estimator:
$\hat{\theta}_{r,\ell_{1}}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;,$
$\implies\hat{\theta}_{r,\ell_{1},0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;.$
then Lemma 4 implies that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
While for the surrogate reward model, note that $\hat{r}^{\star}_{p,0}$ is dense, and $e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$ is intractable, it thus cannot apply $\ell_{0}$-constrained or $\ell_{1}$-regularized estimator to decrease the sample complexity.
Proof of Theorem 11. We set $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$, where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector. For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$. And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank. Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\left(2\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
where $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
Recall that:
$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=2\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
We can have a $\ell_{1}$-regularized estimator for the reward model:
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 5 implies there exists appropriate $\gamma,\delta$, such that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
Now observe the empirical MLE objective of surrogate reward learning:
$\hat{\theta}_{p}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{2})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
where $\tau_{2}:=e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$. Then we can directly apply Lemma 1 and obtain:
$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\Omega\left(\frac{d}{n}\right)\;.$
C.11 Proof of Lemma 5
Lemma 6 (Lemma D.4 of Yao et al. [2025]).
$\mathcal{L}_{\textup{MLE}}(\theta^{\star}+\theta^{\prime})-\mathcal{L}_{\textup{MLE}}(\theta^{\star})-\nabla\mathcal{L}_{\textup{MLE}}(\theta^{\star})^{\top}\theta^{\prime}\geq\Theta(\|\theta^{\prime}\|_{\Sigma_{\mathcal{D}}}^{2})\;,$
for $\forall\theta^{\prime}\in\mathbb{R}^{d}$ s.t. $\theta^{\prime}+\theta^{\star}\in\Theta_{B}$.
We take $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ and $\delta$ are determined in Theorem 3.3 of Yao et al. [2025]. By the definition of the relative $\ell_{1}$-regularized estimator, we have:
$\mathcal{L}_{\textup{MLE}}(\hat{\theta}_{\textup{rel}\ell_{1}})+\gamma\|\hat{\theta}_{\textup{rel}\ell_{1}}-\tau\|_{1}\leq\mathcal{L}_{\textup{MLE}}(\theta^{\star})+\gamma\|\theta^{\star}-\tau\|_{1}$
$\iff$ $\gamma\|\theta^{\star}-\tau\|_{1}-\gamma\|\hat{\theta}_{\textup{rel}\ell_{1}}-\tau\|_{1}\geq\mathcal{L}_{\textup{MLE}}(\hat{\theta}_{\textup{rel}\ell_{1}})-\mathcal{L}_{\textup{MLE}}(\theta^{\star})\;.$
By Lemma 6, we have:
$\mathcal{L}_{\textup{MLE}}(\hat{\theta}_{\textup{rel}\ell_{1}})-\mathcal{L}_{\textup{MLE}}(\theta^{\star})-\nabla\mathcal{L}_{\textup{MLE}}(\theta^{\star})^{\top}(\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star})\geq\Theta(\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2})\;.$
Thus
$\quad\Theta(\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2})\leq\gamma\|\theta^{\star}-\tau\|_{1}-\gamma\|\hat{\theta}_{\textup{rel}\ell_{1}}-\tau\|_{1}-\nabla\mathcal{L}_{\textup{MLE}}(\theta^{\star})^{\top}\left[(\hat{\theta}_{\textup{rel}\ell_{1}}-\tau)-(\theta^{\star}-\tau)\right]$
$\leq\gamma\|\theta^{\star}-\tau\|_{1}-\gamma\|\hat{\theta}_{\textup{rel}\ell_{1}}-\tau\|_{1}+\|\nabla\mathcal{L}_{\textup{MLE}}(\theta^{\star})\|_{\infty}\|\hat{\theta}_{\textup{rel}\ell_{1}}-\tau\|_{1}++\|\nabla\mathcal{L}_{\textup{MLE}}(\theta^{\star})\|_{\infty}\|(\theta^{\star}-\tau)\|_{1}\;,$
where the second inequality is by Hölder’s inequality. Next, we upper bound $\|\nabla\mathcal{L}_{\textup{MLE}}({\theta}^{\star})\|_{\infty}$. As shown in Appendix D.3 of Yao et al. [2025], w.p. $1-\delta$, we have $\|\nabla\mathcal{L}_{\textup{MLE}}({\theta}^{\star})\|_{\infty}\leq\gamma$. Thus, w.p. $1-\delta$, we have:
$\Theta(\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2})$ $\leq\left(\|\nabla\mathcal{L}_{\textup{MLE}}({\theta}^{\star})\|_{\infty}+\gamma\right)\|{\theta}^{\star}-\tau\|_{1}+\left(\|\nabla\mathcal{L}_{\textup{MLE}}({\theta}^{\star})\|_{\infty}-\gamma\right)\|{\hat{\theta}_{\textup{rel}\ell_{1}}-\tau}\|_{1}$
$\leq 2\gamma\|{\theta}^{\star}-\tau\|_{1}\;,$
$\implies\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}$ $=\mathcal{O}(\gamma\|{\theta}^{\star}-\tau\|_{1})\;.$
Note that $\theta^{\star},\tau\in\Theta_{B}$, thus $\|\theta^{\star}-\tau\|_{2}=\mathcal{O}(B)$. Then by Cauchy-Schwartz inequality and the fact that $\|\theta^{\star}-\tau\|_{0}=k$, we have $\|\theta^{\star}-\tau\|_{1}=\mathcal{O}(B\sqrt{k})$, and finally:
$\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
C.12 Omitted Calculations in Sections 3 and B
Calculation of the performance gap. Since $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
$V_{r^{\star}}^{\pi^{\star}}=\underset{y\sim\pi^{\star}}{\mathbb{E}}r(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=\beta\log Z=r^{\star}(y^{\prime})-\beta\log\frac{\pi^{\star}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\;,\;\forall y^{\prime}\in\mathcal{Y}\;.$
Then as motivated by Nika et al. [2024], we can split the performance gap of two-stage RLHF as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
and for DPO we also have:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
Calculation of the underlying “real” objective. When ground-truth reward is non-realizable for the reward model, while the reward model is realizable for the policy model, for a given reward model $r_{\phi}$, the policy model outputs the policy $\pi_{\theta^{\star}(r_{\phi})}$ which satisfies:
$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\,V_{r_{\phi}}^{\pi_{\theta}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\,\underset{y\sim\pi_{\theta}}{\mathbb{E}}r_{\phi}(y)-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\;.$
The solution is given by:
$\pi_{\theta^{\star}(r_{\phi})}(y)=\frac{1}{Z(\phi)}\pi_{{\mathsf{ref}}}(y)\exp\left(\frac{1}{\beta}r_{\phi}(y)\right)\;,$
where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp(r_{\phi}(y)/\beta)$ is the partition function.
The goal of preference-based policy learning is to find a policy $\pi_{\theta}$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. Thus, the reward learning should aim to find $r_{\phi}$ that maximizes:
$V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi_{\theta^{\star}(r_{\phi})}(y)}{\pi_{{\mathsf{ref}}}(y)}\right]$
$=\beta\log Z(\phi)+\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]\;,$
which does not align with maximizing MLE.
Note that
$\nabla_{\phi}\left\{\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]\right\}$ $=\underbrace{\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}[r^{\star}(y)-r_{\phi}(y)]}_{\text{term 1}}-\underbrace{\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)}_{\text{term 2}}\;.$
And we have:
term 1 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r_{\phi}(y)]$
$=\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r^{\star}(y^{\prime})-r_{\phi}(y)+r_{\phi}(y^{\prime})]$ (policy gradient theorem)
$=\frac{1}{2}\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\left[\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)-\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y^{\prime})\right][r^{\star}(y)-r^{\star}(y^{\prime})-r_{\phi}(y)+r_{\phi}(y^{\prime})]\;,$
and
term 2 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)$
$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\beta\nabla_{\phi}\left[\log\pi_{{\mathsf{ref}}}(y)+\log\exp(r_{\phi}(y)/\beta)\right]$
$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\beta\nabla_{\phi}\left[\log\pi_{{\mathsf{ref}}}(y)+\log\exp(r_{\phi}(y)/\beta)-\log Z(\phi)\right]+\beta\nabla_{\phi}\log Z(\phi)$
$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\beta\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)+\beta\nabla_{\phi}\log Z(\phi)$
$=\beta\nabla_{\phi}\log Z(\phi)\;.$ (policy gradient theorem)
By combining them, we obtain Equation 3 and Equation 4.
Note that
$\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+\sigma(r^{\star}(y^{\prime})-r^{\star}(y))\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;,$
and
$\nabla_{q}\left[\sigma(p)\log\sigma(q)+\sigma(-p)\log\sigma(-q)\right]$ $=\sigma(p)\sigma(-q)-\sigma(-p)\sigma(q)$
$=\sigma(p)(1-\sigma(q))-(1-\sigma(p))\sigma(q)$
$=\sigma(p)-\sigma(q)\;,$
we have:
$\nabla_{\phi}\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;,$
which is Equation 5.
To further align the MLE objective with the underlying “real” objective, we can have:
$\nabla_{\phi}\mathcal{L}_{\textup{MLE}}(\phi)\approx-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;,$
and we can assign the value of $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ to the sampling probability $\mu(y,y^{\prime})$. Thus we expect $\mu(y,y^{\prime})\propto\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$. And under the context of surrogate reward learning, we have $\pi_{\theta^{\star}(r_{\phi})}=\pi_{\theta}$ and $r_{\phi}=\hat{r}_{\theta}$, and thus $\mu\propto\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, which is exactly PILAF sampler (see Definition 1).
Calculation of online IPO. For online IPO, let’s observe its objective function:
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ $=\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]^{2}+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y^{\prime})-r_{\theta}(y))-\frac{1}{2}\right]^{2}\;,$
and its gradient is:
$\quad\nabla_{\theta}\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left\{p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))+\frac{1}{2}\right]\right\}\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})\;,$
thus we have:
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
C.13 Omitted Calculations in Section 4
Calculation of the performance gap. Motivated by Nika et al. [2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi^{\star}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$ ($\mathcal{F}\subseteq\mathcal{F}_{\Pi}$)
$=\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
and for DPO we have:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi^{\star}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
Calculation of token-level structure of the optimal solution for DPO. As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
where the $q$ function is determined in a recursive way:
$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
To prove this, we define a $q^{\prime}$ function as:
$q^{\prime}(y_{0})=Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
where $Z:=\log\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$. And then for a $y$ with $y_{N}$ as the terminal token, we have:
$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}$ $=\sum_{t=0}^{N}\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}$
$=q^{\prime}(y_{0})-Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
$=-Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
Note that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, we have:
$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (7)
Then by definition:
$q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
we have:
$\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
which yields:
$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q(s|y_{0\ldots t-1})/\beta)\;.$ (8)
Combining Equations 7 and 8, we complete the calculations.
Appendix D Implementation Details
Codebases. Our codebase is mainly based on MODPO [Zhou et al., 2024] (https://github.com/ZHZisZZ/modpo), Online-RLHF [Dong et al., 2024, Xiong et al., 2024] (https://github.com/RLHFlow/Online-RLHF), Samplers-in-Online-DPO [Shi et al., 2025] (https://github.com/srzer/Samplers-in-Online-DPO). We are committed to releasing the codes.
Datasets. We adopt one common training datasets, PKU-SafeRLHF [Ji et al., 2023] (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF). SFT: We train our initial model on $5$k samples of PKU-SafeRLHF-QA (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-QA). Online training: We use $10$k samples of PKU-SafeRLHF-Prompt (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-prompt) for training, and $2$k samples for evaluation. Offline training: We adopt two preference datasets, PKU-SafeRLHF-safer and PKU-SafeRLHF-better, each composed of $9$k training samples and $2$k evaluation samples, following the practice of Zhou et al. [2024].
Models. Limited by computation resources, our base model is GPT-2-large-774M [Radford et al., 2019] (https://huggingface.co/openai-community/gpt2-large). Our reward model is GPT2-large-harmless model [Yang et al., 2024] (https://huggingface.co/Ray2333/gpt2-large-harmless-reward_model).
Hyper-parameters. The maximum length is set as $256$. The prompt template is “BEGINNING OF CONVERSATION: USER: [prompt] ASSISTANT: [response]”. SFT: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$. Online training: The hyper-parameter setting is based on Dong et al. [2024]. We use a batch size $32$, a learning rate $5\text{e}-7$, and a gradient accumulation step $2$. We train for $3$ iterations, each for $2$ epochs. We set $r_{\textup{margin}}=0.4,1,4$ for verifications of Condition 1, and set $r_{\textup{margin}}=1$ for verifications of Conditions 2, 3 and 4. Offline training: The hyper-parameter setting is based on Zhou et al. [2024]. We use a batch size $4$, a learning rate $1\text{e}-4$, and a gradient accumulation step $2$. We train for $3$ epochs (when training reward model on $9$k data of PKU-SafeRLHF-safer, we train $6$ epochs for higher training accuracy). We haven’t extensively tuned these hyper-parameters.
Computation resources. Our experiments are conducted on NVIDIA RTX A6000. SFT and Online training: We adopt $4$ workers, each taking up $35,000$M of memory, running for $2$-$3$ hours. Offline training: We adopt $1$ worker, which takes up $25,000$M of memory and runs for up to $40$ minutes.

### v2

@@
 In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities.
-We show that RLHF, DPO, or online DPO can outperform one another depending on the type of model mis-specifications.
+We show that RLHF, DPO, or online DPO can outperform one another depending on type of model mis-specifications.
 Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified.
@@
 Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
-[…]
 1 Introduction
@@
 In the policy optimization stage, the base LM is “online” fine-tuned with RL algorithms such as proximal policy optimization (PPO, Schulman et al.
-(2017)), based on $r_{\phi}$ under a Kullback-Leibler (KL) divergence-regularized multi-armed bandit setting.
+(2017)), based on $r_{\phi}$ under a Kullback-Leibler (KL) divergence-regularized bandit setting.
 And the key assumption behind this two-stage pipeline is the realizability of the ground-truth reward.
@@
 Importantly, under the realizability assumption, it is straightforward to derive the equivalence between the ideal performances of RLHF and DPO (Swamy et al., 2025).
-However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., $6$B vs.
-$175$B in Ouyang et al.
-(2022), indicating a clear disparity in representational capacity), when the policy model class is heavily mis-specified due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data.
-Consequently, one should not expect DPO to perform identically to RLHF.
+However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., 66B vs.
+175175B in Ouyang et al.
+(2022), indicating a clear disparity in representational capacity), when the policy model class is heavily restricted due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data.
+These situations are examples of model mis-specification, a common issue in practice due to limitations in model capacity or data.
+Consequently, one should not expect DPO to perform identically to RLHF under model mis-specifications.
 This motivates the central question of our investigation:
 Under what conditions is DPO equivalent, superior, or inferior to RLHF in performance?
-To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($y$ is a response and we omit prompts): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{y\sim\pi}[r^{\star}(y)]-\beta\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)$, where $\pi$ is a distribution over responses, $\pi_{{\mathsf{ref}}}$ is the base model.
+To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$, where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses, and $\pi_{{\mathsf{ref}}}$ is the base model.
 Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
@@
 Sec.
-3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi)\propto-V_{r^{\star}}^{\pi}+\mathcal{O}(\delta^{2})$ Sec.
+3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ Sec.
 3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec.
@@
 We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$.
-Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\mathcal{O}(\sqrt{k\log d/n})$.
-This result indicates that DPO is less sample-efficient than RLHF, leading to inferior performance.
-Finally, we use empirical experiments to corroborate these theoretical findings in Section 5.
+Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$.
+This result indicates that DPO is less data-efficient than RLHF, leading to inferior performance.
+Finally, we conduct numerical experiments to corroborate these theoretical findings in Section 5.
 2 Preliminaries
-[…]
-Multi-armed Bandits and Policies.
-A multi-armed bandit is defined by an action space $\mathcal{Y}$ and a reward function $r:\mathcal{Y}\to\mathbb{R}$.
-A policy $\pi\in\Delta(\mathcal{Y})$ represents a probability distribution over the action space.
-Given the autoregressive mechanism of LM, it is natural to consider the prompt and response together and view the combination as a whole action.
-In this study, our results are presented in the context of multi-armed bandits.
+Notation.
+Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be the sigmoid function, where $\sigma(x)=1/(1+\exp(-x))$.
+For any set $\mathcal{X}$, $\Delta(\mathcal{X})$ represents the set of probability distributions over $\mathcal{X}$.
+$\mathsf{sg}\left(\right)$ is the stopping-gradient operator, where $\nabla_{\theta}[\mathsf{sg}\left(f(\theta)\right)]=\boldsymbol{0}$.
+Let $e_{k}$ be a one-hot vector with $1$ on its $k\textsuperscript{th}$ entry and 0 on other entries.
+We use $f(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}g(\theta)$ to indicate $\nabla_{\theta}f(\theta)=\nabla_{\theta}g(\theta)$.
+Bandits and Policies.
+A bandit is defined by a state space $\mathcal{X}$, an action space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$.
+A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over actions given a state.
+Note that, we sometimes omit the prompt $x$ for simplicity, so that $\pi\in\Delta(\mathcal{Y})$.
 Model class and value function.
-Let =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d_{R}}\}$ denote the reward model class, and $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d_{P}}\}$ denote the policy model class, where $d_{R},d_{P}\in\mathbb{N}$.
+Let $\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d_{R}}\}$ denote the reward model class, and $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d_{P}}\}$ denote the policy model class, where $d_{R},d_{P}\in\mathbb{N}$.
 For a reward function $r$ and policy $\pi$, we define the regularized value function as:
-$V_{r}^{\pi}:=\underset{y\sim\pi}{\mathbb{E}}[r(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;,$
+$V_{r}^{\pi}:=\underset{x\sim\rho}{\mathbb{E}}\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
 where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy.
 Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$.
-A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, where $Z:=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$ is the partition function.
+A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y|x)=\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)/Z(x)$, where $Z(x):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)$ is the partition function.
 The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$.
@@
 Bradley-Terry (BT) model.
-Given an implicit reward oracle $r:\mathcal{Y}\to\mathbb{R}$, Bradley and Terry (1952) assume that human preference distribution $p^{\star}:\mathcal{Y}\times\mathcal{Y}\rightarrow\Delta(\{0,1\})$ satisfies:
-$p^{\star}(y_{1}\succ y_{2})=\sigma\left(r^{\star}(y_{1})-r^{\star}(y_{2})\right)\;.$
-This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2})\;$ by human annotators.
+Given an implicit reward oracle $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, Bradley and Terry (1952) assume that human preference distribution $p^{\star}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\Delta(\{0,1\})$ satisfies:
+$p^{\star}(y_{1}\succ y_{2}|x)=\sigma\left(r^{\star}(x,y_{1})-r^{\star}(x,y_{2})\right)\;.$
+This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2}|x)\;$ by human annotators.
 Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the population MLE objective:
-$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;.$
-In practice, people can have a pre-collected human preference dataset $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where in the $i^{\text{th}}$ sample, $y_{w}^{(i)}\succ y_{l}^{(i)}$ means that $y_{w}^{(i)}$ is preferred over $y_{l}^{(i)}$.
-And the empirical loss is:
-$\mathcal{L}_{\textup{RLHF}}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\;.$
-Then using the reward model $r_{\textup{RLHF}}$, the policy learning stage returns $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\textrm{RLHF}}}^{\pi}$.
+$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))+p^{\star}(y^{\prime}\succ y|x)\log\sigma(r_{\phi}(x,y^{\prime})-r_{\phi}(x,y))\right]\;.$
 Direct approach of DPO.
-By leveraging the surrogate reward $\hat{r}_{\theta}(y):=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
-$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
+By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
+$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
 We also consider an online variant of DPO, where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy.
 Online DPO then minimizes the modified loss:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
-3 Exact Optimization: Fine-grained Representation Gap Induced by Model Mis-specification
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
+3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
 We analyze the behavior of RLHF and DPO in the idealized setting of exact optimization, where both methods have access to infinite preference data and can optimize their respective objectives without statistical or computational error.
@@
 We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes.
-Most prior works (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025) focus on this setting.
+While this assumption is often unrealistic in practice, it serves as a clean baseline and has been the main focus of most prior theoretical analyses (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025).
 Condition 1 (Strong Reward Model, Strong Policy Model).
@@
 In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage.
-Proof deferred to Section C.1.
+Proof deferred to Section B.1.
 Proposition 1.
 Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$.
+Although RLHF and DPO share a same solution, they differ in optimization trajectories and convergence rates.
 Shi et al.
@@
 […]
-with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;$.
+with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
 Remark 1.
-To sample a pairwise data $(y_{1},y_{2})$ from PILAF sampler, we sample $y_{1}\sim\pi^{\mathsf{s1}},y_{2}\sim\pi^{\mathsf{s2}}$, w.p.
-$\alpha_{1}/(\alpha_{1}+\alpha_{2})$ from sampler ➀, and w.p.
-$\alpha_{2}/(\alpha_{1}+\alpha_{2})$ from sampler ➁.
+Given a prompt $x$, we first randomly choose a sampler pair: select sampler ➀ w.p.
+$\alpha_{1}/(\alpha_{1}+\alpha_{2})$ and sampler ➁ otherwise.
+Then sample $y_{1}\sim\pi^{\mathsf{s1}}(\cdot\mid x)$ and $y_{2}\sim\pi^{\mathsf{s2}}(\cdot\mid x)$.
 Theorem 2.
-Given $R_{\max},\delta\in\mathbb{R}_{+}$, s.t.
-$0\leq r^{\star}(y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\delta$, $\forall y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;,$
+Given $R_{\max},\delta\in\mathbb{R}_{+},x\in\mathcal{X}$, s.t.
+$0\leq r^{\star}(x,y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(x,y)-r^{\star}(x,y^{\prime}))-(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))|\leq\delta$, $y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$
+$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(x,y)-r^{\star}(x,y^{\prime})\right)-\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)\right]^{2}\right]\Bigg\}\;,$
 where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t.
-$|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z(\theta):=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$ can be viewed as an adaptive step size.
+$|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
 Remark 2.
-This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value objective up to a second-order deviation.
-However, this deviation can become substantial when the reward function has a large scale or the ground-truth reward is poorly fitted.
+This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value function in prompt level.
+However, the second-order deviation can become substantial when $R_{\max}$ is large, or the ground-truth reward is poorly fitted.
 In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence.
-Proof deferred to Section C.2.
+Proof deferred to Section B.2.
 3.2 Policy Model Mis-specification
@@
 The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization.
-Proof deferred to Section C.3.
+Proof deferred to Section B.3.
 Proposition 3.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
-Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section C.8.
+Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section B.8.
 Proposition 4.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi^{\textup{online}}_{\textup{DPO}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}$ where the online sampler is PILAF sampler (Definition 1).
+Remark 3.
+Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap, an important nuance that, to our knowledge, has been overlooked in prior studies.
 3.3 Reward Model Mis-specification
 We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$.
-This setting is studied for completeness.
 As discussed in Swamy et al.
@@
 Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error.
-Proof deferred to Section C.4.
+Proof deferred to Section B.4.
 Proposition 5.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
+Observation under token-level parameterization.
+To assess the practicality of Condition 3 for auto-regressive language models, we specialize our general bandit model to the token-level parameterization.
+In this setting, the optimal policy admits the closed-form characterization of Rafailov et al.
+(2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Section B.11 for details):
+$\pi^{\star}(y_{t}|x,y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|x,y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|x,y_{0\ldots t-1})}{\beta}\right)\;,$ (1)
+where the $q^{\star}$ function is determined in a recursive way:
+[…]
+This observation shows that while the reward model in RLHF only needs to approximate $r^{\star}$, the policy model in DPO must capture the token-level $q^{\star}$ function, which recursively entangles the reward signal with the base model $\pi_{{\mathsf{ref}}}$.
+As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification and suggesting that the “weak reward, strong policy model” regime may be less common in practice.
 3.4 Double Model Mis-specification
@@
 $r^{\star}\notin\mathcal{F}$, $\pi^{\star}\notin\Pi$.
-To enable a fine-grained comparison between RLHF and DPO under this double mis-specified regime, we introduce the surrogate reward model class induced by the policy class as $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\theta\in\mathbb{R}^{d_{P}},\hat{r}_{\theta}(y)=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)},\forall y\in\mathcal{Y}\}$.
-Pairwise preferences depend only on reward differences, so reward functions $r,r^{\prime}:\mathcal{Y}\to\mathbb{R}$ are equivalent if they differ by a constant.
+To enable a fine-grained comparison between RLHF and DPO under this double mis-specified regime, we introduce the surrogate reward model class induced by the policy class as $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\theta\in\mathbb{R}^{d_{P}},\hat{r}_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{{\mathsf{ref}}}(y|x)},\forall x\in\mathcal{X},y\in\mathcal{Y}\}$.
+Pairwise preferences depend only on reward differences, so reward functions are equivalent if they differ by a constant.
 We compare the expressiveness of the original reward model class $\mathcal{F}$ and the surrogate class $\mathcal{F}_{\Pi}$, modulo constant shifts, and analyze three representative regimes characterizing their relative capacities:
@@
 Note that certain cases involve partially overlapping model classes. However, we do not consider these intermediate regimes for the sake of a principled analysis.
-For Condition 5, we focus on the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies.
+Analysis of the isomorphic case.
+Condition 5 indicates the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies.
 This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage.
-Then motivated by Nika et al.
-(2024), we can split the performance of RLHF and DPO as (see detailed calculations in Section C.12):
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
-In two-stage RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling.
-Therefore, two-stage RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO.
-Proofs deferred to Sections C.5 and C.9.
+In RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling.
+Therefore, RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO.
+Proofs deferred to Sections B.5 and B.9.
 Proposition 6.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
-On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections C.6 and C.7.
+On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections B.6 and B.7.
 Proposition 8.
@@
 $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\textup{RLHF}}}^{\pi}\;,\;\pi_{\textup{DPO}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{\hat{r}_{\textup{DPO}}}^{\pi}\;.$ (2)
-This result implies a general principle: to measure the performance gap, we only need to measure the quality gap between the (surrogate) reward models.
+This result implies a general principle: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$.
+Better reward learning yields higher expected value.
 As revealed in Appendix C of Ouyang et al.
 (2022) and Section 3.3 of Swamy et al.
-(2025), it is uncommon to deploy a reward model with greater scale than the policy model.
-To ensure practical relevance, we focus on the regime where $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
+(2025), it is uncommon to deploy a reward model with a larger scale than the policy model.
+And thus to ensure practical relevance, we focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
 What key property enables a (surrogate) reward model to subsequently help learn good policies?
-As an answer to this question, we can show that in the context of preference-based policy learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference: $\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}$, where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$.
-The detailed calculations and further discussions are deferred to Appendix B.
-In Section 4, we adopt a data-dependent proxy for this notion: the empirical $\ell_{2}$ distance over observed preference pairs (see Definition 2 in Section C.10, also in Zhu et al.
-(2023)).
-Using this metric, we establish a separation result that demonstrates the statistical advantage of explicitly learning reward functions.
-4 Approximate Optimization: Implicit Representation Gap Induced by Statistical Efficiency in Sparse Recovery
-Under more realisitc finite-sample regime, statistical efficiency can also induce performance gap.
-With limited preference data, we are not able to directly compute $r_{\textup{RLHF}}$, $\pi_{\textup{RLHF}}$ and $\pi_{\textup{DPO}}$, and thus obtain weaker solutions due to estimation error.
-This scenario can be viewed as inducing an implicit representation gap, whose effects have been widely discussed in Section 3.
-Motivated by Nika et al.
-(2024) and still assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance of two-stage RLHF as:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}\leq\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{implicit reward representation gap}}\;,$
-and for DPO we have (see detailed calculations in Section C.13):
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $\leq\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
-We now compare the statistical efficiencies of RLHF and DPO, focusing on how their sample complexities differ and influence the implicit reward representation gap.
-We will further improve the analysis of Nika et al.
-(2024); Swamy et al.
-(2025), to prove that when the reward model and the policy model share the same model size, i.e.
-$d_{P}=d_{R}$ and even under realizability condition, two-stage RLHF is still more sample-efficient than DPO.
-Difference in token-level parameterization.
+As an answer to this question, we note that in the context of preference learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference, derived by simple calculations:
+$V_{r_{\phi}}^{\pi_{\theta^{\star}(r_{\phi})}}\textit{ can be measured by }-\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;,$ (3)
+where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$ and we omit prompts for simplicity.
+Detailed calculations and further discussions deferred to Appendix A.
+Using this metric, we can further establish a separation.
+Concluding remarks.
+Although we adopt relatively simple techniques, these results can provide valuable insights for the fundamental differences between RLHF and DPO.
+In the next section, we demonstrate that these insights extend naturally to more practical and realistic scenarios.
+4 Approximate Optimization: Performance Gap Induced by Statistical Efficiency Differences in Reward Learning
+With limited preference data, we are not able to directly compute exact solutions, and thus obtain weaker reward models and policy models due to estimation error.
+This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Condition 4 of Section 3.
+And since we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling.
+Thus by assuming $F\subseteq F_{\Pi}$, we only need to compare the reward model quality measure shown in Equation 3.
+We adopt an empirical proxy for this notion, data-induced semi-norm (details in Definition 2 in Section B.10, see also Zhu et al.
+(2023)): $\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\right]^{2}$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset and we omit prompts from now on.
+Difference in token-level linear parameterization.
+In this section, to rigorously establish a separation, we focus on a specific token-level linear parameterization, which is a special case of the general bandit model; therefore, previous results continue to hold.
 The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value.
-Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta}(y)=\beta\sum_{t=0}^{N}\theta_{t}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and $N=|y|-1$.
+Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta_{r}}(y)=\beta\sum_{t=0}^{n}{\theta_{r,t}}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and n=|y|1n=|y|-1.
 While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al.
 (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
-$\pi_{\theta}(y)=\prod_{t=0}^{N}\pi_{\theta}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{N}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
-Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{N}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$.
-Then the optimal solution for the reward model is $\theta_{t}=\theta^{\star}_{t}$, $\forall t\in\mathbb{Z}_{\geq 0}$.
-And the optimal solution for the policy model is given by Rafailov et al.
-(2024).
-Here we interpret the result in a slightly different way to separate $\pi_{{\mathsf{ref}}}$ and the $q$ function (see detailed calculations in Section C.13):
-$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$ (2)
-where the $q$ function is determined in a recursive way:
-$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
-Benefiting from the token-level $q$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025).
-However, simultaneously, learning the $q$ function sacrifices statistical efficiency due to the need to model the complicated structure.
+$\pi_{\theta_{p}}(y)=\prod_{t=0}^{n}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{n}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
+Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{n}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$, then the optimal solution for the reward model is $\theta_{r,t}^{\star}=\theta^{\star}_{t}$.
+And recall Equation 1, the optimal solution for the policy model is:
+$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
+Benefiting from the token-level $q^{\star}$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025).
+However, simultaneously, learning the $q^{\star}$ function sacrifices statistical efficiency due to the need to model the complicated structure.
 Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
-Dual-token Sparse Prediction (DTSP).
-Under log-linear policy class, the policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
+Dual-token sparse prediction (DTSP) task.
+The policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
 We let $\theta_{r}^{\star}$ denote the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ the optimal solution for surrogate reward learning.
@@
 Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2.
-And even without model mis-specification, we can show that an implicit representation gap can still arise due to statistical inefficiencies.
-Theorem 10 (Efficient Sparse Recovery).
-For DTSP task, and under mild assumptions, there exists an environment, s.t.
-by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(k\log d/n)$ using an $\ell_{0}$-constrained estimator, and can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator.
-Theorem 11 (Separation).
-For DTSP task, and under mild assumptions, there exists an environment, s.t.
-by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using an $\ell_{1}$-regularized estimator, while the estimation error of the policy model is lower bounded by $\Omega(d/n)$.
+And even without explicit model mis-specification, we can establish a separation in reward model qualities due to statistical efficiency differences.
+Theorem 10 (Informal).
+Under token-level linear parameterization and mild assumptions, there exists an environment for DTSP task, s.t.
+by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator, i.e., w.p.
+$1-\delta$,
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r}}(y_{w}^{(i)})-r_{\hat{\theta}_{r}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
+while the estimation error of the DPO model $\hat{\theta}_{p}$ is lower bounded by $\Omega(d/n)$:
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
 Remark 5.
-Benefiting from fixing optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to one-token prediction, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense.
-Pure reward learning can thus reduce the estimation error to $\mathcal{O}(1)$ using $\mathcal{O}\left(k\log d\right)$ samples instead of $\mathcal{O}(d)$ samples.
-Detailed proofs deferred to Section C.10.
-Takeaways.
-This section shows that the policy model is more prone to model mis-specification than the reward model, under a realistic setting, and thus Conditions 1 and 2 are more prevalent than Condition 3.
-And even without model mis-specification, the estimation error can also induce an implicit representation gap.
-From the perspective of sparse recovery, we can see that the surrogate reward model could suffer from severe statistical inefficiency, while the reward model can recover the reward structure very efficiently, even with the same model scale.
+By fixing the optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to a single-token prediction problem, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense.
+Leveraging the results of Yao et al.
+(2025) then yields the separation.
+Formal statement and detailed proof deferred to Section B.10.
+Concluding remarks.
+This section shows that the estimation error can also induce an implicit model mis-specification.
+From the perspective of sparse recovery, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even with the same model scale.
+Although our task construction is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function.
 Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model.
@@
 We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023).
-We first finetune a GPT-2-large-774M model (Radford et al., 2019) on $5$k samples of PKU-SafeRLHF-QA, and obtain the SFT model.
+We first fine-tune a GPT-2-large-774M model (Radford et al., 2019) on 55k samples of PKU-SafeRLHF-QA, and obtain the SFT model.
 We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle.
 All experiments are repeated for $3$ seeds.
-Please see Appendix D for more details.
+Please see Appendix C for more details.
 Implementation details.
@@
 For approximate optimization, we instead compute the empirical BT loss.
-We adopt a pairwise variant of policy gradient instead of PPO to improve training stability: $\mathcal{L}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1}),\hat{r}_{\theta}(y_{2}))\right]^{2}$.
-During deployment, the ground-truth reward will be scaled by a coefficient $r_{\textup{margin}}$.
+We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$.
+During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$.
 Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
@@
 Verifications of Section 4.
-We train offline DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al.
-(2024), and we train on two types of preference: “better” and “safer”, and down-sampled the corresponding training datasets to $1$k-$9$k samples.
+We train DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al.
+(2024).
+We train on two types of preference: “better” and “safer”, and down-sample the corresponding training datasets to 11k-99k samples.
 For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head.
-The models are trained under the same setting, and achieve at least $85\%$ training accuracy.
+The models are trained under the same setting, and all achieve at least $85\%$ training accuracy.
 Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
-Figure 2: Experimental Results for Condition 1. We experiment on different reward scales $0.4,1,4$, and observe that the advantage of RL over online DPO is larger for larger reward scales.
+Figure 2: Experimental Results for Condition 1.
+Experiments with different reward scales $\{0.4,1,4\}$ align with Theorem 2: as the reward scale increases, the second-order deviation in the online DPO objective grows, giving RLHF a clear advantage.
 Figure 3: Experimental Results for Conditions 2, 3 and 4.
-The first two figures align with our claims.
-And the gap in the last figure can be attributed to the mis-specified reward model being too weak.
-Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types, and pure reward learning is shown to be more data-efficient than surrogate reward learning.
-6 Related Works
+The first two plots (Conditions 2 and 3) are consistent with Propositions 3 and 5.
+The gap in the last plot can be attributed to the mis-specified reward model being too weak.
+Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types. Pure reward learning is shown to be more data-efficient than surrogate reward learning.
+6 Related Work
+Reinforcement learning from human feedback (RLHF).
+Seminal contributions that showcased RLHF’s applicability to LLMs include foundational work by Christiano et al.
+(2017), and subsequent research focusing on tasks such as summarization (Stiennon et al., 2020), instruction following (Ouyang et al., 2022), question answering using web-retrieved information (Nakano et al., 2021), and broader AI alignment objectives (Bai et al., 2022).
+Theoretical studies of RLHF include pessimism in policy learning (Zhu et al., 2023), overoptimization (Zhu et al., 2024; Liu et al., 2024c), online RLHF (Xiong et al., 2024; Song et al., 2024), robustness (Mandal et al., 2025), and reward models (Wang et al., 2024; Razin et al., 2025b; Huang et al., 2025; Yao et al., 2025).
+Direct preference optimization (DPO).
+There is a rich literature studying offline (Rafailov et al., 2024; Feng et al., 2024), iterative (Dong et al., 2024; Liu et al., 2024a), and online (Guo et al., 2024; Tajwar et al., 2024; Ding et al., 2024; Shi et al., 2025; Chen et al., 2025; Feng et al., 2025) DPO.
+There are other DPO-style algorithms to directly optimize the policy model from preference signals, such as $\Psi$-PO (Azar et al., 2023), RSO (Liu et al., 2024b), RS-DPO (Khaki et al., 2024), CPO (Xu et al., 2024a), SimPO (Meng et al., 2024), XPO (Xie et al., 2024), VPO (Cen et al., 2024), and OAIF (Guo et al., 2024).
 Performance gap between RLHF and DPO.
@@
 (2024) provided sub-optimality upper bounds for RLHF and DPO when assuming linear reward class and log-linear policy class, with the un-regularized value as performance metric.
-Three cases were studied: ➀ realizable ground-truth reward and exact optimization, ➁ realizable ground-truth reward but approximate optimization, as well as ➂ non-realizable reward and exact optimization.
+Three cases were studied: 1) realizable ground-truth reward and exact optimization, 2) realizable ground-truth reward but approximate optimization, as well as 3) non-realizable reward and exact optimization.
 Let $n$ be the size of the fixed dataset and $d$ be the feature dimension.
-For case ➀, both algorithms have a policy bias due to the un-regularized metric, while RLHF has an additional $\Theta(\sqrt{d/n})$ statistical error and that for DPO is $\Theta(d/(\beta n))$.
-For case ➁, RLHF and DPO both obey a linear convergence to statistical errors and policy biases when using projected gradient descent.
-For case ➂, aside from statistical errors and policy biases, RLHF has an extra approximation error between the ground-truth reward and best achievable reward, while DPO has an extra bias between the optimal regularized policy and the ideal optimal regularized policy.
-Now we focus on comparing with the most relevant prior study, Nika et al.
+For case 1, both algorithms have a policy bias due to the un-regularized metric, while RLHF has an additional $\Theta(\sqrt{d/n})$ statistical error and that for DPO is $\Theta(d/(\beta n))$.
+For case 2, RLHF and DPO both obey a linear convergence to statistical errors and policy biases when using projected gradient descent.
+For case 3, aside from statistical errors and policy biases, RLHF has an extra approximation error between the ground-truth reward and best achievable reward, while DPO has an extra bias between the optimal regularized policy and the ideal optimal regularized policy.
+Here, we focus on comparing with the most relevant prior study, Nika et al.
 (2024).
 First, unlike Nika et al.
-(2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: ➀ it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; ➁ it can help circumvent the unavoidable upper bound of policy bias in the unregularized version.
+(2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable upper bound of policy bias in the unregularized version.
 Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes.
 Third, we improve the statistical analysis of Nika et al.
-(2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\mathcal{O}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
-A more general list of related works is deferred to Appendix A.
+(2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\tilde{\mathcal{O}}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
 7 Conclusion
 This paper provides a fine-grained analysis of the performance gap between two-stage and direct approaches to preference-based policy learning.
-We theoretically demonstrate a dichotomy of RLHF and DPO under different mis-specification scenarios, and further reveal a implicit representation gap induced by statistical efficiency.
+We theoretically demonstrate a dichotomy of RLHF and DPO under different mis-specification scenarios, and further reveal an implicit representation gap induced by statistical efficiency.
 Our claims are supported by empirical experiments on LMs.
 It is also important to acknowledge our limitations.
-➀ While we identify a limitation of training reward models based on BT model, we do not provide a theoretically grounded and practically effective alternative.
-➁ Due to computational constraints, our experiments are limited to small-scale models.
+1) While we identify a limitation of training reward models based on BT model, we do not provide a theoretically grounded and practically effective alternative.
+2) Due to computational constraints, our experiments are limited to small-scale models.
 We hope our insights can motivate the community to further investigate these directions.
 Acknowledgement
-SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P.
+SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P.
 Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship.
-RZ and MF acknowledge the support of NSF TRIPODS II DMS-2023166.
-The work of MF was supported in part by awards NSF CCF 2212261 and NSF CCF 2312775.
+RZ acknowledges the support of NSF TRIPODS II DMS-2023166.
 RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
 Appendix
-[…]
-Appendix A Supplementary Related Works
-Reinforcement learning from human feedback (RLHF).
-Seminal contributions that showcased RLHF’s applicability to LLMs include foundational work by Christiano et al.
-[2017], and subsequent research focusing on tasks such as summarization [Stiennon et al., 2020], instruction following [Ouyang et al., 2022], question answering using web-retrieved information [Nakano et al., 2021], and broader AI alignment objectives [Bai et al., 2022].
-Theoretical studies of RLHF include pessimism in policy learning [Zhu et al., 2023], overoptimization [Zhu et al., 2024, Liu et al., 2024c], online RLHF [Xiong et al., 2024, Song et al., 2024], robustness [Mandal et al., 2025], and reward models [Wang et al., 2024, Razin et al., 2025b, Huang et al., 2025, Yao et al., 2025].
-Direct preference optimization (DPO).
-There is a rich literature studying offline [Rafailov et al., 2024, Feng et al., 2024], iterative [Dong et al., 2024, Liu et al., 2024a], and online [Guo et al., 2024, Tajwar et al., 2024, Ding et al., 2024, Shi et al., 2025, Chen et al., 2025, Feng et al., 2025] DPO.
-There are other DPO-style algorithms to directly optimize the policy model from preference signals, such as $\Psi$-PO [Azar et al., 2023], RSO [Liu et al., 2024b], RS-DPO [Khaki et al., 2024], CPO [Xu et al., 2024a], SimPO [Meng et al., 2024], XPO [Xie et al., 2024], VPO [Cen et al., 2024], and OAIF [Guo et al., 2024].
-Appendix B Bonus: How Can We Better Model Reward From Preference Signals?
+Appendix A Bonus: How Can We Better Model Reward From Preference Signals?
 As motivated by Equation 2, a reward model $r_{\phi}$ can be mapped to a policy via:
@@
 $r_{\phi^{\star}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmin}}\;\underbrace{-\beta\log Z(\phi)-\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:\mathcal{L}_{\textup{new}}(\phi)}\;.$
-Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section C.12):
+Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section B.11):
 $\nabla_{\phi}\mathcal{L}_{\textup{new}}(\phi)=-\frac{1}{2}\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;,$ (4)
@@
 And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$.
-If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it is a process reward model.
-However, it is worth noting that the model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2.
-Therefore, when faced with representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
-To alleviate this issue, the remedy could be to learn the preference with alternative modeling approaches to circumvent the BT model setting, which has already be revealed in the successes of Sun et al.
+If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it can provide token-level reward information.
+However, it is worth noting that model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2.
+Therefore, when faced with a representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
+To alleviate this issue, we could learn the preference with alternative modeling approaches to circumvent the BT model setting, which has already shown success in Sun et al.
 [2025], Calandriello et al.
 [2024].
-For example, we can look into the training objective of online IPO [Calandriello et al., 2024, Zhou et al., 2025] (see detailed calculations in Section C.12):
+For example, we can look into the training objective of online IPO [Calandriello et al., 2024, Zhou et al., 2025] (see detailed calculations in Section B.11):
 $\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y_{1},y_{2})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))-\frac{p^{\star}(y_{1}\succ y_{2})-p^{\star}(y_{2}\succ y_{1})}{2}\right]^{2}\;,$
-where $\rho(\theta)$ is an online sampling distribution, and it thus can optimize a $\ell_{2}$ distance in an online way.
+where $\rho(\theta)$ is an online sampling distribution, and it thus can optimize an $\ell_{2}$ distance in an online way.
 The classification model deployed in Sun et al.
 [2025] is also promising.
-We leave this interesting direction for future exploration, and now come back to two-stage RLHF and DPO.
-Appendix C Omitted Proofs
-C.1 Proof of Proposition 1
+We leave this interesting direction for future exploration.
+Appendix B Omitted Proofs
+Note that in this section, we omit all prompts without loss of generality.
+For the constructive proof, we can set the number of states to $1$; for the other proofs, we can simply sum over different prompts to extend them to the general case.
+B.1 Proof of Proposition 1
 Since $r^{\star}\in\mathcal{F}$, RLHF exactly recovers $r^{\star}$ during reward learning.
@@
 Hence, $\pi_{\textup{DPO}}=\pi^{\star}$, which achieves the maximum of $V_{r^{\star}}^{\pi}$ over $\Pi$.
-C.2 Proof of Theorem 2
+B.2 Proof of Theorem 2
 By Taylor’s expansion, we have that:
-$\nabla_{\theta}\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ $=-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]$
+$\quad\nabla_{\theta}\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta})$
+$=-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]$
 $\quad-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime\prime}(\xi_{y,y^{\prime}})\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]^{2}\;,$
@@
 then the formula can be rewritten as:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{2\mathsf{sg}\left(Z(\theta)\right)}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left(1+\epsilon_{y,y^{\prime}}\right)\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;,$
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta})$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{2\mathsf{sg}\left(Z_{\theta}\right)}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left(1+\epsilon_{y,y^{\prime}}\right)\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;,$
 where
-$|\epsilon_{y,y^{\prime}}|$ $=\left|\frac{\sigma^{\prime\prime}(\xi_{y,y^{\prime}})}{\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))}\right|\cdot|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|$
-$\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime}(R_{\max}+\delta)}\;,$
+$|\epsilon_{y,y^{\prime}}|$ $=\left|\frac{\sigma^{\prime\prime}(\xi_{y,y^{\prime}})}{\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))}\right|\cdot|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime}(R_{\max}+\delta)}\;,$
 and
-$Z(\theta):=\sum_{y,y^{\prime}\in\mathcal{Y}}\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;.$
+$Z_{\theta}:=\sum_{y,y^{\prime}\in\mathcal{Y}}\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;.$
 Note that:
-$\nabla_{\theta}V_{r^{\star}}^{\pi_{\theta}}$ $=\nabla_{\theta}\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(y)-\hat{r}_{\theta}(y)]$
+$\quad\nabla_{\theta}\left[\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\right]$ (7)
+$=\nabla_{\theta}\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(y)-\hat{r}_{\theta}(y)]$
 $=\underset{y\sim\pi_{\theta}}{\mathbb{E}}\nabla_{\theta}\log\pi_{\theta}(y)[r^{\star}(y)-\hat{r}_{\theta}(y)]$ (policy gradient theorem)
@@
 thus
-$V_{r^{\star}}^{\pi_{\theta}}$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}-\frac{1}{4\beta}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;.$
+$\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}-\frac{1}{4\beta}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;.$
 Therefore we have:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;.$
-C.3 Proof of Proposition 3
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\right]$
+$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\Bigg\}\;.$
+B.3 Proof of Proposition 3
 Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$.
@@
 Now consider the DPO objective, which relies on pairwise preference probabilities and directly optimizes over the policy class:
-$\mathcal{L}_{\textup{DPO}}(\theta)$ $=-\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma\left(\beta\,\theta^{\top}(\psi(a)-\psi(a^{\prime})\right)\right]$
+$\mathcal{L}_{\textup{DPO}}(\pi_{\theta})$ $=-\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma\left(\beta\,\theta^{\top}(\psi(a)-\psi(a^{\prime})\right)\right]$
 $=-\tfrac{1}{2}\log\sigma(\beta\Delta^{\top}\theta)-\tfrac{1}{2}\log\sigma(-\beta\Delta^{\top}\theta)-\log\sigma(\tfrac{1}{2}\beta\Delta^{\top}\theta)-\log\sigma(-\tfrac{1}{2}\beta\Delta^{\top}\theta)\;,$
@@
 Thus, unlike RLHF, the DPO solution remains fixed at uniform distribution, independent of the reward magnitude $r$ and the regularization parameter $\beta$, and fails to suppress the sub-optimal action $a_{3}$ even when $\beta$ is sufficiently small.
-C.4 Proof of Proposition 5
+B.4 Proof of Proposition 5
 Since $r^{\star}\notin\mathcal{F}$, RLHF recovers an approximation $\hat{r}\in\mathcal{F}$ via reward learning.
@@
 Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V_{\Pi}^{\star}$.
-For the strict gap, consider a multi-armed bandit setting analogous to Section C.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
+For the strict gap, consider a multi-armed bandit setting analogous to Section B.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
 $r=r^{\star}(a_{1})=r^{\star}(a_{2})\geq r^{\star}(a_{3})=0\;.$
@@
 The key difference from the earlier construction lies in the choice of model classes.
-We define: the linear reward class =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d}\}$ by $r_{\phi}(a):=\phi^{\top}\psi(a)$, and the policy class $\Pi=\Delta(\mathcal{Y})$ with reference policy $\pi_{{\mathsf{ref}}}=\textrm{Unif}(\mathcal{Y})$.
+We define: the linear reward class $\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d}\}$ by $r_{\phi}(a):=\phi^{\top}\psi(a)$, and the policy class $\Pi=\Delta(\mathcal{Y})$ with reference policy $\pi_{{\mathsf{ref}}}=\textrm{Unif}(\mathcal{Y})$.
 This setup satisfies Condition 3 because $r^{\star}\notin\mathcal{F}$: for any $\phi$, the constraint on $\psi$ implies $r_{\phi}(a_{3})=\tfrac{1}{2}(r_{\phi}(a_{1})+r_{\phi}(a_{2}))$ so $r_{\phi}(a_{3})=r$ whenever $r_{\phi}(a_{1})=r_{\phi}(a_{2})=r$, contradicting the ground-truth reward $r^{\star}(a_{3})=0$.
@@
 where $\Delta:=\psi(a_{1})-\psi(a_{2})$.
-This expression is minimized at $\Delta^{\top}\phi=0$, which implies $r_{\phi}(a_{1})=r_{\phi}(a_{2})$ and $r_{\phi}(a_{3})=r_{\phi}(a_{1})$, i.e., the learned reward is constant: $r_{\textup{RLHF}}(a)=C$ for all $a\in\mathcal{Y}$.
+This expression is minimized maximized at $\Delta^{\top}\phi=0$, which implies $r_{\phi}(a_{1})=r_{\phi}(a_{2})$ and $r_{\phi}(a_{3})=r_{\phi}(a_{1})$, i.e., the learned reward is constant: $r_{\textup{RLHF}}(a)=C$ for all $a\in\mathcal{Y}$.
 The policy learning stage then solves:
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}\;.$
-C.5 Proof of Proposition 6
+B.5 Proof of Proposition 6
 By definition, the reward learned by RLHF and the surrogate reward learned by DPO are obtained by solving the following population objectives:
@@
 $\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}\quad\text{and hence}\quad V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
-C.6 Proof of Proposition 8
+B.6 Proof of Proposition 8
 Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
 We first construct an environment satisfying Condition 6 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
-Consider the same setup as in Section C.4, but define the policy class as $\Pi=\Delta(\mathcal{Y})\setminus\{\pi^{\star}\}$, where $\pi^{\star}$ is the optimal policy under $r^{\star}$.
+Consider the same setup as in Section B.4, but define the policy class as $\Pi=\Delta(\mathcal{Y})\setminus\{\pi^{\star}\}$, where $\pi^{\star}$ is the optimal policy under $r^{\star}$.
 This ensures that $\pi^{\star}\notin\Pi$, while $\mathcal{F}\subset\mathcal{F}_{\Pi}$, satisfying Condition 6.
-As shown in Section C.4, RLHF learns a constant reward model and returns the uniform policy $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
+As shown in Section B.4, RLHF learns a constant reward model and returns the uniform policy $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
 In contrast, DPO directly optimizes policy parameters from preference data and can converge to a policy arbitrarily close to $\pi^{\star}$, which lies on the boundary of $\Pi$.
@@
 The corresponding surrogate reward class is $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a),\;\theta\in\mathbb{R}^{2}\}$.
-We now define a strictly smaller reward model class =$\mathcal{F}=\{\hat{r}_{\theta_{R}}\}$ where
+We now define a strictly smaller reward model class $\mathcal{F}=\{\hat{r}_{\theta_{R}}\}$ where
 $\theta_{R}=\begin{bmatrix}1\\ -1\end{bmatrix}$.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=\pi_{\theta_{R}}(a_{1})+\pi_{\theta_{R}}(a_{2})-\beta\,\mathrm{KL}(\pi_{\theta_{R}}\|\pi_{{\mathsf{ref}}})\approx 0.729\;.$
-In contrast, DPO learns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, as shown in Section C.3. Its value is:
+In contrast, DPO learns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, as shown in Section B.3. Its value is:
 $V_{r^{\star}}^{\pi_{\textup{DPO}}}=\frac{2}{3}\;.$
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}\;.$
-C.7 Proof of Proposition 9
+B.7 Proof of Proposition 9
 Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
@@
 Then $\mathcal{F}_{\Pi}\subset\mathcal{F}$, and thus Condition 7 holds.
-From Section C.3, we know that DPO learns a constant reward model under this feature structure and returns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
+From Section B.3, we know that DPO learns a constant reward model under this feature structure and returns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
 RLHF, on the other hand, optimizes the MLE objective over the larger class $\mathcal{F}$ and selects $\bar{r}$, which achieves a higher likelihood than any function in $\mathcal{F}_{\Pi}$.
@@
 $V_{r^{\star}}^{\pi_{\theta}}$ $=\pi_{\theta}(a_{1})+\pi_{\theta}(a_{2})-\beta\,\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)$
-[…]
+$=\frac{e^{x/2}+e^{-x/2}}{Z(x)}-\beta\left[\frac{e^{x/2}}{Z(x)}\log\left(\frac{e^{x/2}}{Z(x)}\right)+\frac{e^{-x/2}}{Z(x)}\log\left(\frac{e^{-x/2}}{Z(x)}\right)+\frac{1}{Z(x)}\log\left(\frac{1}{Z(x)}\right)\right]+\text{(constant)}\;,$
 where $x:=\theta^{\top}\begin{bmatrix}1\\ -1\end{bmatrix}$ and $Z(x):=e^{x/2}+e^{-x/2}+1$.
@@
 demonstrating that a more expressive reward model class may lead RLHF to overfitting in the presence of a constrained policy class, resulting in inferior performance compared to DPO.
-C.8 Numerical Proof of Proposition 4
+B.8 Numerical Proof of Proposition 4
+Since the exact solution for online DPO is hard to compute, we didn’t find elegant proofs for these two propositions. They are examined correct numerically.
 By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
@@
 We find that both DPO and online DPO will converge to the same sub-optimal solution, while RL can obtain an optimal solution.
-C.9 Numerical Proof of Proposition 7
+B.9 Numerical Proof of Proposition 7
 By Proposition 6, we have $\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}$.
 Now we only need to construct an environment under Condition 5, such that online DPO can outperform offline DPO.
-We can borrow the whole setting in Section C.8, while resetting the ground-truth reward as:
+We can borrow the whole setting in Section B.8, while resetting the ground-truth reward as:
 $r^{\star}(a_{1})=24\;,\;r^{\star}(a_{2})=12\;,\;r^{\star}(a_{3})=0\;.$
@@
 Figure 5: Numerically Computed Curves of Gradient Functions and Value Functions.
-C.10 Proofs of Theorems 10 and 11
+B.10 Formal Statement of Theorem 10 and Proof
+B.10.1 Preliminaries of single-token prediction
+Before proceeding, we first prepare some ingredients for the single-token prediction task.
+Basic setting.
 Recall that to train a (surrogate) reward model, people first collect a dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$.
@@
 $y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p.
-$\sigma(r^{\star}(y_{1})-r^{\star}(y_{2})$, where $r^{\star}(y)=(\theta^{\star})^{\top}\psi(y)$, $\theta^{\star}\in\mathbb{R}_{+}$ is the ground-truth reward vector, $\psi(y)$ is the feature vector satisfying $\|\psi(y)\|_{2}\leq L$, and $L\in\mathbb{R}_{+}$.
+$\sigma(r^{\star}(y_{1})-r^{\star}(y_{2}))$, where $r^{\star}(y)=(\theta^{\star})^{\top}\psi(y)$, $\theta^{\star}\in\mathbb{R}_{+}$ is the ground-truth reward vector, $\psi(y)$ is the feature vector satisfying $\|\psi(y)\|_{2}\leq L$, and $L\in\mathbb{R}_{+}$.
 The MLE estimator is defined as:
-$\hat{\theta}_{\textup{MLE}}=\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\theta^{\top}(\psi(y_{w})-\psi(y_{l})))\;,$ (6)
+$\hat{\theta}_{\textup{MLE}}\in\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\theta^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$ (8)
 where $\Theta_{B}=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $B\in\mathbb{R}_{+}$.
 And we assume $\theta^{\star}\in\Theta_{B}$.
-The data-dependent performance measure is the data-induced semi-norm [Zhu et al., 2023], which is defined as:
+The empirical performance measure is the data-induced semi-norm (see, e.g., [Zhu et al., 2023]), defined as:
 Definition 2 (Data-induced semi-norm).
@@
 $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}\;.$
-Remark 5.
-Note that $\hat{\theta}_{1}$ has been fixed as $e_{1}$ in our context, we thus only need to measure $\hat{\theta}_{0}$.
-Before proceeding, we first prepare some ingredients. It is worth noting that those lemmas below only work for the one-token scenario, and we will adopt them in the dual-token prediction task later.
+And we assume $\Sigma_{\mathcal{D}}$ to be non-singular.
+Note that the lemmas below only work for the single-token scenario, and we will adopt them in the dual-token prediction task later.
+The results quoted below from [Yao et al., 2025] follow directly from a long line of work on compressed sensing and sparse recovery based on restricted isometry (or restricted eigenvalue) properties [Candes et al., 2006], recast for the preference learning setting.
 Lemma 1 (Theorem 1.a of Shah et al. [2015]).
-For a sample size $n\geq c_{1}\text{tr($\Sigma_{\mathcal{D}}$)}$, any estimator $\hat{\theta}$ based on $n$ samples has a lower bound as:
+For a sample size $n\geq c_{1}\text{tr($\Sigma_{\mathcal{D}}^{-1}$)}$, any estimator $\hat{\theta}$ based on $n$ samples has a lower bound as:
 $\underset{\theta^{\star}\in\Theta_{B}}{\sup}\mathbb{E}\left[\|\hat{\theta}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}\right]=\Omega\left(\frac{d}{n}\right)\;.$
 Remark 6.
-Here $c_{1}$ is a numerical constant.
-This lemma is to establish a information-theoretical lower bound of one-token reward learning.
-Therefore, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ defined in Equation 6 is lower bounded.
-Lemma 2 (Lemma 3.1 of Zhu et al. [2023]).
-With probability at least $1-\delta$, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ has an upper bound:
+Here $c_{1}$ is a constant independent of data. This lemma is to establish an information-theoretical lower bound for single-token reward learning.
+Lemma 2 (Lemma 3.1 of Zhu et al. [2023]; see also Shah et al. [2015]).
+W.p. at least $1-\delta$, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ has an upper bound:
 $\|\hat{\theta}_{\textup{MLE}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;.$
-Definition 3 ($\ell_{0}$-constrained estimator).
-$\hat{\theta}_{\ell_{0}}^{k}\in\underset{\theta\in\Theta_{B},\|\theta\|_{0}\leq k}{\operatorname{argmin}}\;\mathcal{L}_{\textup{MLE}}(\theta)\;.$
-Lemma 3 (Theorem 3.2 of Yao et al. [2025]).
-Consider $\|\theta^{\star}\|_{0}=k$ and $k\leq d/2$, then with probability at least $1-\delta$, the $\ell_{0}$-constrained estimator $\hat{\theta}_{\ell_{0}}^{k}$ has an upper bound:
-$\|\hat{\theta}_{\ell_{0}}^{k}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
 Definition 3 ($\ell_{1}$-regularized estimator).
@@
 Lemma 3 (Theorem 3.3 of Yao et al. [2025]).
-Consider $\|\theta^{\star}\|_{0}=k$, then with probability at least $1-\delta$, the $\ell_{1}$-regularized estimator $\hat{\theta}_{\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$ has an upper bound:
-$\|\hat{\theta}_{\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\sigma)}{n}}\right)\;.$
+Consider $\|\theta^{\star}\|_{0}=k$, then w.p.
+at least $1-\delta$, the $\ell_{1}$-regularized estimator $\hat{\theta}_{\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\sqrt{\frac{\log(d)+\log(1/\delta)}{n}}\right)$ has an upper bound:
+$\|\hat{\theta}_{\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
 Definition 4 (Relative $\ell_{1}$-regularized estimator).
@@
 Lemma 4 (Generalized version of Lemma 3).
-Consider $\tau\in\Theta_{B}$, $\|\theta^{\star}-\tau\|_{0}=k$, then with probability at least $1-\delta$, the relative $\ell_{1}$-regularized estimator $\hat{\theta}_{\textup{rel}\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\frac{\log 2d+\log(1/\delta)}{n}\right)$ has an upper bound:
-$\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
-Proof of this lemma is deferred to Section C.11.
-Proof of Theorem 10.
-We first prepare the settings and assumptions.
-For the feature mapping, we set $\psi(y,\omega)=\psi(\omega)$.
-For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs.
-Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$.
-And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank, and $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$.
-Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:
-$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})=\left(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
-where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector and $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
-$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
-which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
-Recall that:
-$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
-$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})+C_{4}\;,$
-we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
-Then observe the empirical MLE objective of pure reward learning:
-$\hat{\theta}_{r}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
-$\implies\hat{\theta}_{r,0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
-and the empirical MLE objective of surrogate reward learning:
-$\hat{\theta}_{p}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
-$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
-We can thus directly apply Lemma 2 to obtain that w.p. $1-\delta$, we have:
-$\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;,$
-and
-$\|\hat{\theta}_{r,0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;.$
-Note that $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$, we can have an $\ell_{0}$-constrained estimator for the reward model:
-$\hat{\theta}_{r,\ell_{0}}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k,\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
-$\implies\hat{\theta}_{r,\ell_{0},0}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
-then Lemma 3 implies that w.p. $1-\delta$,
-$\|\hat{\theta}_{r,\ell_{0},0}^{k}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
-And we also have an $\ell_{1}$-regularized estimator:
-$\hat{\theta}_{r,\ell_{1}}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;,$
-$\implies\hat{\theta}_{r,\ell_{1},0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;.$
-then Lemma 4 implies that w.p. $1-\delta$,
-$\|\hat{\theta}_{r,\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
-While for the surrogate reward model, note that $\hat{r}^{\star}_{p,0}$ is dense, and $e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$ is intractable, it thus cannot apply $\ell_{0}$-constrained or $\ell_{1}$-regularized estimator to decrease the sample complexity.
-Proof of Theorem 11.
-We set $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$, where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector.
-For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs.
-Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$.
-And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank.
-Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:
-$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\left(2\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
-where $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
-$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
-which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
+Consider $\tau\in\Theta_{B}$, $\|\theta^{\star}-\tau\|_{0}=k$, then w.p.
+at least $1-\delta$, the relative $\ell_{1}$-regularized estimator $\hat{\theta}_{\textup{rel}\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\frac{\log(d)+\log(1/\delta)}{n}\right)$ has an upper bound:
+$\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
+Proof of this lemma is given in Section B.10.4.
+B.10.2 Formal statement of Theorem 10
+Assumption 11 (Task configuration).
+Recall that in DTSP task, we have $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)$, where $\psi(y),\psi(y,\omega),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$.
+We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(y)\|_{2}\leq L$, and $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$.
+Assumption 12 (Preference data collection).
+For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs.
+Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)}$ and $y_{l}^{(i)}=\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}$.
+And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))^{\top}$ is non-singular, and $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$.
+Theorem 13 (Formal separation theorem).
+Under token-level linear parameterization and Assumptions 11 and 12, there exists an environment for DTSP task, s.t.
+by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator:
+$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+e_{1}+\textbf{r}_{\textup{dense}}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
+i.e., w.p. $1-\delta$,
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{w}^{(i)})-r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
+while the estimation error of any estimator for the DPO model $\hat{\theta}_{p}$ is lower bounded by $\Omega(d/n)$:
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+B.10.3 Proof of Theorem 13
+Let $\pi_{{\mathsf{ref}}}(\cdot|y)$ be identical for all $y$, then we have
+$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(y)+C_{5}\;,$
+for $\forall y$, where $C_{5}\in\mathbb{R}$ is an offset.
 Recall that:
@@
 $(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
-we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=2\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
+we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ and $\theta^{\star}_{p,0}=\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
 We can have a $\ell_{1}$-regularized estimator for the reward model:
-$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
-$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
-where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 5 implies there exists appropriate $\gamma,\delta$, such that w.p. $1-\delta$,
-$\|\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
-Now observe the empirical MLE objective of surrogate reward learning:
-$\hat{\theta}_{p}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
-$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{2})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
-where $\tau_{2}:=e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$. Then we can directly apply Lemma 1 and obtain:
-$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\Omega\left(\frac{d}{n}\right)\;.$
-C.11 Proof of Lemma 5
+$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)})-r_{\theta}(\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
+$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
+where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 4 implies there exists appropriate $\gamma$, such that w.p. $1-\delta$,
+$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
+and thus w.p. $1-\delta$,
+$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{w}^{(i)})-r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\theta_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\theta_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
+$=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
+Note that
+$\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
+then Lemma 1 implies that $\exists N=\mathcal{O}(d)$, for $\forall n\geq N$ and any estimator $\hat{\theta}_{p}$, we have
+$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+Now observe the data-induced semi-norm of surrogate reward learning:
+$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}\;.$
+And thus there exists an environment for DTSP, s.t.
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+B.10.4 Proof of Lemma 4
 Lemma 5 (Lemma D.4 of Yao et al. [2025]).
@@
 for $\forall\theta^{\prime}\in\mathbb{R}^{d}$ s.t. $\theta^{\prime}+\theta^{\star}\in\Theta_{B}$.
-We take $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ and $\delta$ are determined in Theorem 3.3 of Yao et al.
+We take $\gamma=\Theta\left(\sqrt{\frac{\log(d)+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ is determined in Theorem 3.3 of Yao et al.
 [2025].
@@
 $\implies\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}$ $=\mathcal{O}(\gamma\|{\theta}^{\star}-\tau\|_{1})\;.$
-Note that $\theta^{\star},\tau\in\Theta_{B}$, thus $\|\theta^{\star}-\tau\|_{2}=\mathcal{O}(B)$.
-Then by Cauchy-Schwartz inequality and the fact that $\|\theta^{\star}-\tau\|_{0}=k$, we have $\|\theta^{\star}-\tau\|_{1}=\mathcal{O}(B\sqrt{k})$, and finally:
-$\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
-C.12 Omitted Calculations in Sections 3 and B
-Calculation of the performance gap. Since $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
-$V_{r^{\star}}^{\pi^{\star}}=\underset{y\sim\pi^{\star}}{\mathbb{E}}r(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=\beta\log Z=r^{\star}(y^{\prime})-\beta\log\frac{\pi^{\star}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\;,\;\forall y^{\prime}\in\mathcal{Y}\;.$
-Then as motivated by Nika et al. [2024], we can split the performance gap of two-stage RLHF as:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
-and for DPO we also have:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
+Note that $\theta^{\star},\tau\in\Theta_{B}$, thus $\|\theta^{\star}-\tau\|_{2}=\mathcal{O}(1)$.
+Then by Cauchy-Schwartz inequality and the fact that $\|\theta^{\star}-\tau\|_{0}=k$, we have $\|\theta^{\star}-\tau\|_{1}=\mathcal{O}(\sqrt{k})$, and finally:
+$\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
+B.11 Omitted Calculations
+Calculation of token-level structure of the optimal solution for DPO. As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
+$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
+$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\star}(y_{0})-\beta\log Z}{\beta}\right)\;,$
+where $Z:=\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$, and the $q^{\star}$ function is determined in a recursive way:
+[…]
+To prove this, we define a $q^{\prime}$ function as:
+$q^{\prime}(y_{0})=\beta\log Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;.$
+For the initial token, by definition we have:
+$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\prime}(y_{0})-\beta\log Z}{\beta}\right)\;.$ (9)
+And then for a $y$ with $y_{N}$ as the terminal token, we have:
+$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}$ $=\sum_{t=0}^{N}\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}$
+$=q^{\prime}(y_{0})-\beta\log Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
+$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
+Note that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, we have:
+$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=-\beta\log Z+r^{\star}(y)\;,$
+thus
+$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (10)
+Then by definition:
+$q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
+we have:
+$\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=\pi^{\star}(y_{t}|y_{0\ldots t-1})\;,$ (11)
+and thus
+$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
+which yields:
+$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (12)
+Combining Equations 9, 11, 10 and 12, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
 Calculation of the underlying “real” objective.
@@
 And we have:
-term 1 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r_{\phi}(y)]$
+term 1
+$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r_{\phi}(y)]$
 $=\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r^{\star}(y^{\prime})-r_{\phi}(y)+r_{\phi}(y^{\prime})]$ (policy gradient theorem)
@@
 and
-term 2 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)$
+term 2
+$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)$
 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\beta\nabla_{\phi}\left[\log\pi_{{\mathsf{ref}}}(y)+\log\exp(r_{\phi}(y)/\beta)\right]$
@@
 Thus we expect $\mu(y,y^{\prime})\;\propto\;\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$.
-And under the context of surrogate reward learning, we have $\pi_{\theta^{\star}(r_{\phi})}=\pi_{\theta}$ and $r_{\phi}=\hat{r}_{\theta}$, and thus $\mu\propto\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, which is exactly PILAF sampler (see Definition 1).
+And under the context of DPO, we have $\pi_{\theta^{\star}(r_{\phi})}=\pi_{\theta}$ and $r_{\phi}=\hat{r}_{\theta}$, and thus $\mu\;\propto\;\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, which is exactly PILAF sampler.
 Calculation of online IPO. For online IPO, let’s observe its objective function:
-$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ $=\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]^{2}+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y^{\prime})-r_{\theta}(y))-\frac{1}{2}\right]^{2}\;,$
+$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$ $=\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]^{2}+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y^{\prime})-r_{\theta}(y))-\frac{1}{2}\right]^{2}\;,$
 and its gradient is:
-$\quad\nabla_{\theta}\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$
-$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left\{p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))+\frac{1}{2}\right]\right\}\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})$
-$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})\;,$
+$\quad\nabla_{\theta}\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$
+$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left\{p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))+\frac{1}{2}\right]\right\}\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime}))$
+$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime}))\;,$
 thus we have:
-$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
-C.13 Omitted Calculations in Section 4
-Calculation of the performance gap. Motivated by Nika et al. [2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
+$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
+Below we provide two additional results for reference.
+Exact Optimization: calculation of the policy performance gap. Since the optimal policy satisfies $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
+$V_{r^{\star}}^{\pi^{\star}}=\underset{y\sim\pi^{\star}}{\mathbb{E}}r(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=\beta\log Z=r^{\star}(y^{\prime})-\beta\log\frac{\pi^{\star}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\;,\;\forall y^{\prime}\in\mathcal{Y}\;.$
+Then as motivated by Nika et al. [2024], we can split the performance gap of two-stage RLHF as:
+$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
+and for DPO we also have:
+$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
+Approximate Optimization: calculation of the reward representation gap.
+Motivated by Nika et al.
+[2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
 $V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi^{\star}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
@@
 $\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
-$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
-Calculation of token-level structure of the optimal solution for DPO. As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
-$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
-where the $q$ function is determined in a recursive way:
-$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
-To prove this, we define a $q^{\prime}$ function as:
-$q^{\prime}(y_{0})=Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
-where $Z:=\log\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$. And then for a $y$ with $y_{N}$ as the terminal token, we have:
-$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}$ $=\sum_{t=0}^{N}\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}$
-$=q^{\prime}(y_{0})-Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
-$=-Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
-Note that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, we have:
-$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (7)
-Then by definition:
-$q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
-we have:
-$\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
-which yields:
-$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q(s|y_{0\ldots t-1})/\beta)\;.$ (8)
-Combining Equations 7 and 8, we complete the calculations.
-Appendix D Implementation Details
+$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
+Appendix C Implementation Details
 Codebases.
 Our codebase is mainly based on MODPO [Zhou et al., 2024] (https://github.com/ZHZisZZ/modpo), Online-RLHF [Dong et al., 2024, Xiong et al., 2024] (https://github.com/RLHFlow/Online-RLHF), Samplers-in-Online-DPO [Shi et al., 2025] (https://github.com/srzer/Samplers-in-Online-DPO).
-We are committed to releasing the codes.
+Our code is released at https://github.com/srzer/Gap-in-Preference-Learning.
 Datasets.
-We adopt one common training datasets, PKU-SafeRLHF [Ji et al., 2023] (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF).
-SFT: We train our initial model on $5$k samples of PKU-SafeRLHF-QA (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-QA).
-Online training: We use $10$k samples of PKU-SafeRLHF-Prompt (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-prompt) for training, and $2$k samples for evaluation.
-Offline training: We adopt two preference datasets, PKU-SafeRLHF-safer and PKU-SafeRLHF-better, each composed of $9$k training samples and $2$k evaluation samples, following the practice of Zhou et al.
+We adopt one common training dataset, PKU-SafeRLHF [Ji et al., 2023] (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF).
+SFT: We train our initial model on 55k samples of PKU-SafeRLHF-QA (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-QA).
+Online training: We use 1010k samples of PKU-SafeRLHF-Prompt (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-prompt) for training, and 22k samples for evaluation.
+Offline training: We adopt two preference datasets, PKU-SafeRLHF-safer and PKU-SafeRLHF-better, each composed of 99k training samples and 22k evaluation samples, following the practice of Zhou et al.
 [2024].
@@
 We use a batch size $4$, a learning rate $1\text{e}-4$, and a gradient accumulation step $2$.
-We train for $3$ epochs (when training reward model on $9$k data of PKU-SafeRLHF-safer, we train $6$ epochs for higher training accuracy).
+We train for $3$ epochs (when training reward model on 99k data of PKU-SafeRLHF-safer, we train $6$ epochs for higher training accuracy).
 We haven’t extensively tuned these hyper-parameters.
@@
 Our experiments are conducted on NVIDIA RTX A6000.
-SFT and Online training: We adopt $4$ workers, each taking up $35,000$M of memory, running for $2$-$3$ hours.
-Offline training: We adopt $1$ worker, which takes up $25,000$M of memory and runs for up to $40$ minutes.
+SFT and Online training: We adopt $4$ workers, each taking up 35,00035, [^0] M of memory, running for 22-33 hours.
+Offline training: We adopt $1$ worker, which takes up 25,00025, [^0] M of memory and runs for up to $40$ minutes.

### v3

@@
 Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified.
-In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model—highlighting a statistical advantage of two-stage learning.
+In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model, highlighting a statistical advantage of two-stage learning.
 Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
@@
 (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily restricted due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data.
-These situations are examples of model mis-specification, a common issue in practice due to limitations in model capacity or data.
-Consequently, one should not expect DPO to perform identically to RLHF under model mis-specifications.
+These situations are examples of representation gap, a common issue in practice due to limitations in model capacity or data.
+Consequently, one should not expect DPO to perform identically to RLHF.
 This motivates the central question of our investigation:
 Under what conditions is DPO equivalent, superior, or inferior to RLHF in performance?
-To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$, where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses, and $\pi_{{\mathsf{ref}}}$ is the base model.
+To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response):
+$V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
+where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses given prompts, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy.
 Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
-Our contributions. We study the performance differences between two-stage RLHF and DPO under a representation gap, from an optimization perspective. Our contributions are listed as follows:
-Sec.
-3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ Sec.
-3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec.
-3.2 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ Sec.
-3.4 It depends on
+Our contributions. We study the performance differences between RLHF and DPO, from optimization and statistic perspectives. Our contributions are listed as follows:
+$V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ It depends on
 the qualities of
 $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$.
-[…]
+$V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}$$V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$No Model Mis-specificationPolicy Model Mis-specificationReward Model Mis-specificationIsomorphic Model Mis-specification
 Figure 1: Main results on performance gap induced by model mis-specification scenarios.
-$\bullet$ When assuming exact optimization, i.e., optimization with infinite data, we study the fine-grained representation gap under different settings of reward and policy class mis-specifications in Section 3.
+$\bullet$ When assuming exact optimization, i.e., optimization with infinite data, we remove the widely-used realizability assumption, and study the fine-grained performance gap under different settings of model mis-specifications in Section 3.
 Main results are visualized in Figure 1.
-No model mis-specification: We show that the RLHF and DPO policies both achieve the performance of $\pi^{\star}$, and online DPO can further close the gap between optimization paths.
-Policy model mis-specification: We show that the RLHF policy is still optimal under the model class, while the DPO policy can be sub-optimal, and online DPO cannot bridge the gap.
-Reward model mis-specification: We show that the DPO policy is still optimal, while the RLHF policy can be sub-optimal due to learning based on a sub-optimal reward model.
-Double model mis-specification: When policy and reward model classes are isomorphic, then they should have identical performance, while online DPO can outperform both of them.
+No model mis-specification: RLHF and DPO policies both achieve the performance of $\pi^{\star}$, and online DPO can further close the gap between optimization paths.
+Policy model mis-specification: RLHF policy is still optimal under the model class, while the DPO policy can be sub-optimal, and online DPO cannot bridge the gap.
+Reward model mis-specification: DPO policy is still optimal, while the RLHF policy can be sub-optimal due to learning based on a sub-optimal reward model.
+Double model mis-specification: When policy and reward model classes are isomorphic, they should have identical performance, while online DPO can outperform both of them.
 Otherwise, there is no consistent performance gap, and the comparison result depends on the qualities of (surrogate) reward models.
-We also give a preliminary guide for reward learning under mis-specifications.
-$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the implicit representation gap incurred by statistical efficiencies in Section 4.
-We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$.
-Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$.
+$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the performance gap incurred by statistical efficiencies in Section 4.
+We show that DPO training can distort the intrinsic structure of the reward, and demonstrate this effect under a linear parameterization.
+We then construct a simple task where the ground-truth reward feature has dimension $d$ and sparsity $k$, and the total number of samples is $n$.
+Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, reducing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$.
 This result indicates that DPO is less data-efficient than RLHF, leading to inferior performance.
-Finally, we conduct numerical experiments to corroborate these theoretical findings in Section 5.
+Finally, we conduct controlled experiments to corroborate these theoretical findings and connect existing empirical studies with our theoretical claims in Section 5.
 2 Preliminaries
@@
 Let $e_{k}$ be a one-hot vector with $1$ on its $k\textsuperscript{th}$ entry and 0 on other entries.
+For any vector $x$, let $x_{k}$ be its $k\textsuperscript{th}$ entry.
 We use $f(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}g(\theta)$ to indicate $\nabla_{\theta}f(\theta)=\nabla_{\theta}g(\theta)$.
-Bandits and Policies.
-A bandit is defined by a state space $\mathcal{X}$, an action space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$.
-A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over actions given a state.
+We define the difference operator as $\Delta_{x_{1},x_{2}}(f)=f(x_{1})-f_{1}(x_{2})$.
+Policy space.
+There is a prompt space $\mathcal{X}$, a response space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$.
+A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over responses given a prompt.
 Note that, we sometimes omit the prompt $x$ for simplicity, so that $\pi\in\Delta(\mathcal{Y})$.
@@
 For a reward function $r$ and policy $\pi$, we define the regularized value function as:
-$V_{r}^{\pi}:=\underset{x\sim\rho}{\mathbb{E}}\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
-where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy.
+$V_{r}^{\pi(\cdot|x)}$ $:=\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
+$V_{r}^{\pi}$ $:=\underset{x\sim\rho}{\mathbb{E}}V_{r}^{\pi(\cdot|x)}\;,$
+where $\beta>0$ is the regularization coefficient, $\rho\in\Delta(\mathcal{X})$ is a pre-fixed distribution over prompts, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy.
 Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$.
@@
 The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$.
-We define the oracle value as $V^{\star}_{\Pi}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
+We define the oracle value as $V^{\Pi}_{r^{\star}}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
 Bradley-Terry (BT) model.
@@
 This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2}|x)\;$ by human annotators.
-Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the population MLE objective:
-$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))+p^{\star}(y^{\prime}\succ y|x)\log\sigma(r_{\phi}(x,y^{\prime})-r_{\phi}(x,y))\right]\;.$
+Empirical preference dataset.
+In practice, people first collect a pair dataset $\mathcal{D}^{\dagger}=\{x^{(i)},y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$.
+Following BT model, $y_{1}^{(i)}$ is preferred over $y_{2}^{(i)}$ given prompt $x^{(i)}$, (i.e.
+$y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p.
+$p^{\star}(y_{1}^{(i)}\succ y_{2}^{(i)}|x^{(i)})$.
+Two-stage approach of RLHF. RLHF proceeds in two stages. First, the reward learning stage finds a reward model $r_{\textup{RLHF}}\in\mathcal{F}$ by maximizing the MLE objective:
+$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma(r_{\phi}(x,y_{1})-r_{\phi}(x,y_{2}))\;.$
+And for approximate optimization, $r_{\text{RLHF}}$ is estimated from a finite human preference dataset.
+Then using the reward model $r_{\textup{RLHF}}$, the policy learning stage returns $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\textrm{RLHF}}}^{\pi}$.
 Direct approach of DPO.
-By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
-$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
-We also consider an online variant of DPO, where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy.
-Online DPO then minimizes the modified loss:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
+By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns policy from preference:
+$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma\left(\hat{r}_{\theta}(x,y_{1})-\hat{r}_{\theta}(x,y_{2})\right)\;.$
+For approximate optimization, $\pi_{\text{DPO}}$ is estimated from an empirical preference dataset.
+We also consider an online variant of DPO (Xiong et al., 2024), where the pairwise data are sampled from a distribution $\pi^{\mathsf{s}}$ which could depend on the current policy.
+It then minimizes the modified loss:
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\sum_{\{y_{1},y_{2}\}=\{y,y^{\prime}\}}p^{\star}(y_{1}\succ y_{2}|x)\log\sigma\left(\hat{r}_{\theta}(x,y_{1})-\hat{r}_{\theta}(x,y_{2})\right)\;.$
 3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
 We analyze the behavior of RLHF and DPO in the idealized setting of exact optimization, where both methods have access to infinite preference data and can optimize their respective objectives without statistical or computational error.
-Our focus is on the performance gap induced by the representation gap – that is, the difference between the true optimal policy and the best policy each method can produce, as determined by the expressiveness of the reward and policy model classes.
 Recall that $r_{\textup{RLHF}}\in\mathcal{F}$ is the solution computed by exact optimization of reward learning, $\pi_{\textup{RLHF}}\in\Pi$ is the solution computed by exact optimization of policy learning given $r_{\textup{RLHF}}$, and $\pi_{\textup{DPO}}\in\Pi$ is the solution computed by exact optimization of DPO.
-3.1 No Model Mis-specification
+We can bound the sub-optimality of each algorithm using the mis-specification error (see calculations in Appendix B.12), but in this section our focus is on the performance gap induced by model mis-specification, that is, the difference between the best policy each method can produce, as determined by the expressiveness of the reward and policy model classes.
+No Model Mis-specification.
 We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes.
@@
 In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage.
-Proof deferred to Section B.1.
+Proof deferred to Appendix B.1.
 Proposition 1.
-Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$.
+Under Condition 1, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}$.
 Although RLHF and DPO share a same solution, they differ in optimization trajectories and convergence rates.
@@
 […]
-with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
+with ratio $\alpha_{1}=1,\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
 Remark 1.
@@
 $0\leq r^{\star}(x,y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(x,y)-r^{\star}(x,y^{\prime}))-(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))|\leq\delta$, $y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
-$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$
-$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(x,y)-r^{\star}(x,y^{\prime})\right)-\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)\right]^{2}\right]\Bigg\}\;,$
+$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}\Bigg\{-V_{r}^{\pi(\cdot|x)}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\Delta^{2}_{(x,y),(x,y^{\prime})}(r^{\star}-\hat{r}_{\theta})\right]\Bigg\}\;,$
 where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t.
-$|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
+$|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$, $\Delta(\cdot)$ is the difference operator as defined in Section 2, and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
 Remark 2.
@@
 In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence.
-Proof deferred to Section B.2.
-3.2 Policy Model Mis-specification
+Proof deferred to Appendix B.2.
+Policy Model Mis-specification.
 We now examine the setting where the ground-truth reward function is realizable ($r^{\star}\in\mathcal{F}$), but the optimal policy is non-realizable by the policy class ($\pi^{\star}\notin\Pi$).
@@
 The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization.
-Proof deferred to Section B.3.
+Proof deferred to Appendix B.3.
 Proposition 3.
-Under Condition 2, $V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and there exists an environment s.t.
+Under Condition 2, $V^{\Pi}_{r^{\star}}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$, and there exists an environment s.t.
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}>V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
-Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section B.8.
+Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Appendix B.8.
 Proposition 4.
@@
 Remark 3.
-Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap, an important nuance that, to our knowledge, has been overlooked in prior studies.
-3.3 Reward Model Mis-specification
+Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap.
+This is an important nuance that has been overlooked in prior studies.
+Reward Model Mis-specification.
 We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$.
@@
 Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error.
-Proof deferred to Section B.4.
+Proof deferred to Appendix B.4.
 Proposition 5.
-Under Condition 3, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$, and there exists an environment s.t.
+Under Condition 3, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}$, and there exists an environment s.t.
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
 Observation under token-level parameterization.
-To assess the practicality of Condition 3 for auto-regressive language models, we specialize our general bandit model to the token-level parameterization.
+To assess the practicality of Conditions 2 and 3 for auto-regressive language models, we can specialize our general policy class to the token-level parameterization.
 In this setting, the optimal policy admits the closed-form characterization of Rafailov et al.
-(2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Section B.11 for details):
+(2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Appendix B.10 for details):
 $\pi^{\star}$ $(y_{t}|x,y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|x,y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|x,y_{0\ldots t-1})}{\beta}\right)\;,$ (1)
 where the $q^{\star}$ function is determined in a recursive way:
-[…]
+$q^{\star}(y_{t}|x,y_{0\ldots t-1})=\begin{cases}\beta\log\mathbb{E}_{s\sim\pi_{\textup{ref}}(\cdot|x,y_{0\ldots t})}\exp\!\big(q^{\star}(s|x,y_{0\ldots t})/\beta\big)\;\;&\text{if $y_{t}$ is not terminal,}\\ r^{\star}(x,y_{0\ldots t})\;\;\begin{subarray}{c}\end{subarray}&\text{if $y_{t}$ is terminal}.\end{cases}$
+Here $\mathcal{V}$ is the vocabulary, and $s\in\mathcal{V}$ is the token.
 This observation shows that while the reward model in RLHF only needs to approximate $r^{\star}$, the policy model in DPO must capture the token-level $q^{\star}$ function, which recursively entangles the reward signal with the base model $\pi_{{\mathsf{ref}}}$.
-As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification and suggesting that the “weak reward, strong policy model” regime may be less common in practice.
-3.4 Double Model Mis-specification
-We now consider the most challenging setting, where neither the ground-truth reward function nor the optimal policy is realizable by their respective model classes.
+As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification than the reward model of the same scale.
+Therefore, in practice, people sometimes deploy small reward model and can still obtain acceptable performance, while the deployed policy model cannot be too small.
+Double Model Mis-specification. We now consider the most challenging setting, where neither the ground-truth reward function nor the optimal policy is realizable by their respective model classes.
 Condition 4 (Weak Reward Model, Weak Policy Model).
@@
 Therefore, RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO.
-Proofs deferred to Sections B.5 and B.9.
+Proofs deferred to Appendix B.5 and  B.9.
 Proposition 6.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
-On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections B.6 and B.7.
+On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Appendix B.6 and  B.7.
 Proposition 8.
@@
 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
+Understanding the role of reward quality.
 Though there is no consistent performance gap between RLHF and DPO in certain settings, revisiting the framework can reveal a structural parallel: RLHF can yield the best policy given the learned reward model $r_{\textup{RLHF}}$, and the DPO policy is directly the optimal one given the surrogate reward model $\hat{r}_{\textup{DPO}}$.
-And online DPO serves to enhance the quality of $\hat{r}_{\textup{DPO}}$ (Xiong et al., 2024).
+And online DPO/RLHF (Xiong et al., 2024; Li et al., 2025) serves to enhance the quality of $\hat{r}_{\textup{DPO}}$.
 Formally,
 $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\textup{RLHF}}}^{\pi}\;,\;\pi_{\textup{DPO}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{\hat{r}_{\textup{DPO}}}^{\pi}\;.$ (2)
-This result implies a general principle: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$.
+This result implies a general principle holding for all conditions in this section: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$.
 Better reward learning yields higher expected value.
-As revealed in Appendix C of Ouyang et al.
-(2022) and Section 3.3 of Swamy et al.
-(2025), it is uncommon to deploy a reward model with a larger scale than the policy model.
-And thus to ensure practical relevance, we focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
-What key property enables a (surrogate) reward model to subsequently help learn good policies?
-As an answer to this question, we note that in the context of preference learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference, derived by simple calculations:
-$V_{r_{\phi}}^{\pi_{\theta^{\star}(r_{\phi})}}\textit{ can be measured by }-\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;,$ (3)
-where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$ and we omit prompts for simplicity.
-Detailed calculations and further discussions deferred to Appendix A.
-Using this metric, we can further establish a separation.
+We will look into reward model qualities in next section, and defer an algorithmic discussion to Appendix A.
 Concluding remarks.
 Although we adopt relatively simple techniques, these results can provide valuable insights for the fundamental differences between RLHF and DPO.
-In the next section, we demonstrate that these insights extend naturally to more practical and realistic scenarios.
+Later we will demonstrate that these insights extend naturally to more practical and realistic scenarios.
 4 Approximate Optimization: Performance Gap Induced by Statistical Efficiency Differences in Reward Learning
 With limited preference data, we are not able to directly compute exact solutions, and thus obtain weaker reward models and policy models due to estimation error.
-This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Condition 4 of Section 3.
-And since we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling.
-Thus by assuming $F\subseteq F_{\Pi}$, we only need to compare the reward model quality measure shown in Equation 3.
-We adopt an empirical proxy for this notion, data-induced semi-norm (details in Definition 2 in Section B.10, see also Zhu et al.
-(2023)): $\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\right]^{2}$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset and we omit prompts from now on.
-Difference in token-level linear parameterization.
-In this section, to rigorously establish a separation, we focus on a specific token-level linear parameterization, which is a special case of the general bandit model; therefore, previous results continue to hold.
+This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Section 3.
+And because we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling.
+So our main focus will be to compare the qualities of reward model and DPO model, and we omit prompts from now on.
+Notably, for linear bandit model studied in previous studies (Xiong et al., 2024; Zhu et al., 2023; Yao et al., 2025), where the ground truth reward is $r^{\star}(y)=\beta(\theta^{\star})^{\top}\psi(y)$, the reward model is $r_{\theta_{r}}(y)=\beta\theta^{\top}\psi(y)$, and the policy model is $\pi_{\theta_{p}}(y)\propto\pi_{{\mathsf{ref}}}(y)\exp(\theta^{\top}\psi(y))$, we have that $\theta_{r}$ and $\theta_{p}$ share the same optimal solution $\theta^{\star}$, and there is no difference in sample complexity.
+To rigorously establish a separation, we focus on the more realistic token-level linear parameterization, which is a special case of our general policy class; therefore, results in Section 3 continue to hold.
+Reward model.
 The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value.
-Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta_{r}}(y)=\beta\sum_{t=0}^{n}{\theta_{r,t}}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and n=|y|1n=|y|-1.
-While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al.
+Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.:
+$r_{\theta_{r}}(y)=\beta\sum_{t=0}^{|y|-1}\theta_{r,t}^{\top}\psi(y_{0\ldots t})\;,$
+where $\theta_{r,t},\psi(y_{0\ldots t})\in\mathbb{R}^{d}$.
+Policy model.
+While for the policy model, one needs to go through softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al.
 (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
-$\pi_{\theta_{p}}(y)=\prod_{t=0}^{n}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{n}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
-Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{n}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$, then the optimal solution for the reward model is $\theta_{r,t}^{\star}=\theta^{\star}_{t}$.
-And recall Equation 1, the optimal solution for the policy model is:
-$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
-Benefiting from the token-level $q^{\star}$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025).
-However, simultaneously, learning the $q^{\star}$ function sacrifices statistical efficiency due to the need to model the complicated structure.
-Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
+$\quad\pi_{\theta_{p}}(y)=\prod_{t=0}^{|y|-1}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{|y|-1}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s\in\mathcal{V}}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;,$
+where $\theta_{p,t}\in\mathbb{R}^{d}$, $\mathcal{V}$ is the vocabulary, and the surrogate reward model is $\hat{r}_{\theta_{p}}(y)=\beta\sum_{t=0}^{|y|-1}\log\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})$.
+Let $\theta_{r}^{\star}$ be the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ be the optimal solution for DPO.
+Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{|y|-1}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$.
+Then the optimal solution for the reward model is
+$\theta_{r,t}^{\star}=\theta^{\star}_{t}+C\;,$ (3)
+where $C$ is a constant offset, and recall Equation 1, the optimal solution for the policy model is:
+$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;,$
+and thus
+$\left(\theta_{p,t}^{\star}\right)^{\top}\psi(y_{0\ldots t})=q^{\star}(y_{t}|y_{0\ldots t-1})/\beta+C^{\prime}=\left(\theta_{t}^{\star}\right)^{\top}\psi(y_{0\ldots t})+v^{\star}(y_{t}|y_{0\ldots t-1})/\beta+C^{\prime\prime}\;,$ (4)
+where $v^{\star}$ is the soft value function (see detailed derivation in Appendix B.10), and $C^{\prime},C^{\prime\prime}$ are constant offsets.
+Curse of value function.
+Comparing Equation 3 and Equation 4, we can see that the main distinction lies in whether to learn the value function.
+While it is widely recognized that learning $q^{\star}$$v^{\star}$ function can provide fine-grained process information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025), this benefit comes with a significant statistical cost: estimating $q^{\star}/v^{\star}$ generally incurs a higher sample complexity because the induced structure is considerably richer.
+Notably, even under a linear assumption for $q^{\star}/v^{\star}$, the value function couples the reward with the base model, so some good properties of the underlying reward (e.g., sparsity) are not preserved.
+Next, we will present a minimalist example to rigorously illustrate the statistical gap between RLHF and DPO.
+We adopt an empirical proxy, data-induced semi-norm (details in Definition 2 in Appendix B.11, see also Zhu et al.
+(2023)): $\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-r_{\phi})$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset.
 Dual-token sparse prediction (DTSP) task.
-The policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
-We let $\theta_{r}^{\star}$ denote the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ the optimal solution for surrogate reward learning.
-Note that for the second token, $\theta_{r}^{\star}$ and $\theta_{p}^{\star}$ share the same optimal solution:
-$(\theta_{r,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{1}\;,\;(\theta_{p,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{2}\;,$
-where $C_{1},C_{2}\in\mathbb{R}$ are offsets. And for the first token $y$, there is a distinction:
-$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
-$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
-where $\textbf{r}_{\textup{sparse}}$ gets entangled with $\pi_{{\mathsf{ref}}}$ in $\theta_{p,0}^{\star}$.
-Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2.
-And even without explicit model mis-specification, we can establish a separation in reward model qualities due to statistical efficiency differences.
-Theorem 10 (Informal).
+Let $\mathcal{V}$ be the vocabulary, and $\mathcal{Y}=\mathcal{V}^{2}$.
+The policy model is required to sequentially output two tokens $a,b$, and the ground-truth reward is: $r^{\star}(a,b)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\beta e_{1}^{\top}\psi(a,b)\;,$ where $a,b\in\mathcal{V}$, $\psi(a),\psi(a,b)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
+Theorem 10 (Informal reward separation).
 Under token-level linear parameterization and mild assumptions, there exists an environment for DTSP task, s.t.
@@
 $1-\delta$,
-$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r}}(y_{w}^{(i)})-r_{\hat{\theta}_{r}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
+$\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-r_{\hat{\theta}_{r}})=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
 while the estimation error of the DPO model $\hat{\theta}_{p}$ is lower bounded by $\Omega(d/n)$:
-$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+$\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-\hat{r}_{\hat{\theta}_{p}})=\Omega\left(\frac{d}{n}\right)\;.$
 Remark 5.
@@
 (2025) then yields the separation.
-Formal statement and detailed proof deferred to Section B.10.
+Formal statement and detailed proof deferred to Appendix B.11.
+Theorem 11 (Informal sub-optimality separation).
+Under token-level linear parameterization and mild assumptions, there exists an environment for DTSP task, s.t. we can have a separation on the sub-optimality of RLHF and DPO:
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $=\tilde{\mathcal{O}}\left(\sqrt[4]{\frac{k\log d}{n}}\cdot\sqrt{\Lambda_{1}}\right)\;,$
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $=\Omega\left(\frac{d}{n}\cdot\Lambda_{2}\right)\;,$
+where $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\hat{\theta}_{r}}}^{\pi}$, $\pi_{\textup{DPO}}=\pi_{\hat{\theta}_{p}}$, and $\Lambda_{1},\Lambda_{2}$ are geometric quantities of data.
+Formal statement and detailed proof deferred to Appendix B.11.5.
 Concluding remarks.
-This section shows that the estimation error can also induce an implicit model mis-specification.
-From the perspective of sparse recovery, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even with the same model scale.
-Although our task construction is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function.
-Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model.
-This further explains why two-stage RLHF is empirically observed to outperform DPO (Ivison et al., 2024; Xu et al., 2024b).
+In this section, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even under the same model scale.
+Although our parameterization is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function.
+For general model class beyond linear model, Equation 1 still holds.
+And hence, compared with reward model, the policy model always faces a more complex target, making it more vulnerable to model mis-specification and sample inefficiency.
+Additionally, to prevent policy model mis-specification (discussed in Section 3), $d_{P}$ is often required to be larger than $d_{R}$, which further increases sample complexity.
+In the belief that real-world rewards are often simple or sparse (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the DPO model.
+This further explains why two-stage RLHF empirically outperforms DPO in practice (Ivison et al., 2024; Xu et al., 2024b).
 5 Experimental Verifications
+5.1 Controlled experimental verifications
 Experiment setup.
-We now verify our analysis in practical settings.
+We now verify our analysis in practical and controlled settings.
 We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023).
@@
 Please see Appendix C for more details.
-Implementation details.
-For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset.
-For approximate optimization, we instead compute the empirical BT loss.
-We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$.
-During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$.
-Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
 Verifications of Section 3.
@@
 Results are shown in Figures 2 and 3.
+The empirical findings align closely with our theoretical predictions:
+  • Figure 2 (Condition 1) aligns with Propositions 1 and 2: increasing the reward scale amplifies the second-order deviation in online DPO’s objective, causing larger deviation from the RLHF optimum, as our theory predicts.
+  • Figure 3 (left, Condition 2) confirms Propositions 3 and 4: with a realizable reward model but restricted policy class, RLHF outperforms DPO.
+  • Figure 3 (middle, Condition 3) confirms Proposition 5: with a mis-specified reward model but realizable policy class, DPO outperforms RLHF.
+  • Figure 3 (right, Condition 4) exhibits behavior consistent with our double model mis-specification analysis: relative performance can depend on the comparative expressive power of $\mathcal{F}$ versus $\mathcal{F}_{\Pi}$.
+In our setup, the reward model is less expressive, leading RLHF to underperform.
 Verifications of Section 4.
@@
 Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
+This corroborates our theoretical separation result in Theorem 10: pure reward learning is statistically more sample-efficient than the surrogate reward learning performed by DPO.
 Figure 2: Experimental Results for Condition 1.
@@
 Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types. Pure reward learning is shown to be more data-efficient than surrogate reward learning.
+5.2 Connection with exisiting empirical studies.
+Although our experiments are intentionally small-scale, our conclusions are supported by several recent large-scale empirical studies:
+  • Table 1 of Razin et al.
+(2025b) show that DPO is worse than RLHF when learning the reward, supporting our claim in Section 4.
+Their models are Gemma-2-2B-IT, Qwen series, and Llama series, and the datasets are UltraFeedback and RewardMATH.
+  • Figure 4 of Swamy et al. (2025) supports our observation in Section 3 that policy model requires larger scale to prevent mis-specification.
+  • Figure 6 of Swamy et al. (2025) supports our Proposition 4 that online DPO cannot improve DPO much when the policy model is mis-specified. Their model is Pythia-1.4B.
+  • Table 6 of Xu et al.
+(2024b) demonstrates that RLHF is stronger than DPO with a perfect reward signal, supporting our Proposition 3.
+Their models are Llama series, and benchmarks are SafeRLHF and APPS.
+  • Table 5 of Ivison et al.
+(2024) demonstrates that RLHF can beat DPO under a good reward model, supporting our Proposition 3 and double-model-mis-specification analysis.
+Their models are Tulu2 series.
+  • Figure 2 of Rafailov et al.
+(2023) demonstrates that DPO can beat PPO under a weak reward model, supporting our Proposition 5 and double-model-mis-specification analysis.
+Their reward model is initialized from GPT2-large model and trained for only $3$ epochs on the datasets, and is thus relatively weak.
+These large-scale studies support the qualitative trends predicted in our theoretical analysis.
 6 Related Work
@@
 First, unlike Nika et al.
-(2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable upper bound of policy bias in the unregularized version.
-Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes.
+(2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable policy bias in the unregularized version.
+Second, we provide an analysis of model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n$ is large enough, and we are not limited to linear model classes.
 Third, we improve the statistical analysis of Nika et al.
 (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\tilde{\mathcal{O}}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
+Concurrent works.
+We also acknowledge two concurrent works (Razin et al., 2026; Gopalan et al., 2026), which both show that DPO is not an idealized estimator for reward modeling.
+We differ in approaches: Razin et al.
+(2026) study learning dynamics, and Gopalan et al.
+(2026) propose a geometric characterization.
 7 Conclusion
@@
 Acknowledgement
-SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P.
+SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P.
 Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship.
-RZ acknowledges the support of NSF TRIPODS II DMS-2023166.
+RZ and MF acknowledge the support of NSF TRIPODS II DMS-2023166.
+The work of MF was supported in part by awards NSF CCF 2212261 and NSF CCF 2312775.
 RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
@@
 Appendix A Bonus: How Can We Better Model Reward From Preference Signals?
+In this section, we study the following relevant question:
+What key property enables a (surrogate) reward model to subsequently help learn good policies?
 As motivated by Equation 2, a reward model $r_{\phi}$ can be mapped to a policy via:
-$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\phi}}^{\pi}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;\underset{y\sim\pi}{\mathbb{E}}[r_{\phi}(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;.$
-If $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, this solution further admits the closed form $\pi_{\theta^{\star}(r_{\phi})}(y)=\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)/Z(\phi)$, where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)$ is the partition function.
-If the goal is to output a policy that performs well under the ground-truth reward $r^{\star}$, then reward learning should aim to find a model $r_{\phi}$ such that the resulting policy $\pi_{\theta^{\star}(r_{\phi})}$ maximizes the underlying “real” objective:
-$r_{\phi^{\star}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmin}}\;\underbrace{-\beta\log Z(\phi)-\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:\mathcal{L}_{\textup{new}}(\phi)}\;.$
-Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section B.11):
+$\pi_{\theta^{\star}(r_{\phi})}:=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\phi}}^{\pi}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;\underset{y\sim\pi}{\mathbb{E}}[r_{\phi}(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;,$
+and the reward model can further be projected to a surrogate reward model:
+$\hat{r}_{\theta^{\star}(r_{\phi})}(y):=\beta\log\pi_{\theta^{\star}(r)}(y)/\pi_{{\mathsf{ref}}}(y)\;.$
+As revealed in Appendix C of Ouyang et al.
+[2022] and Section 3.3 of Swamy et al.
+[2025], it is uncommon to deploy a reward model with a larger scale than the policy model.
+And thus to ensure practical relevance, we will focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$.
+Since $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, the solution further admits the closed form $\pi_{\theta^{\star}(r_{\phi})}(y)=\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)/Z(\phi)$, where $Z(\phi):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp\left(r_{\phi}(y)/\beta\right)$ is the partition function, and $\hat{r}_{\theta^{\star}(r_{\phi})}=r_{\phi}$ module constant shift.
+If the goal is to output a policy that performs well under the ground-truth reward $r^{\star}$, then reward learning should aim to find a model $r_{\phi}$ such that the resulting policy $\pi_{\theta^{\star}(r_{\phi})}$ optimizing the underlying “real” objective:
+$V_{r^{\star}}^{\pi_{\theta^{\star}(r_{\phi})}}$ $=\underbrace{\beta\log Z(\phi)+\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}[r^{\star}(y)-r_{\phi}(y)]}_{=:-\mathcal{L}_{\textup{new}}(\phi)}\;.$
+Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Appendix B.12):
 $\nabla_{\phi}\mathcal{L}_{\textup{new}}(\phi)=-\frac{1}{2}\underset{y,y^{\prime}\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;,$ (5)
-which corresponds to the gradient of an $\ell_{2}$ distance of pairwise difference:
+which corresponds to the gradient of an $\ell_{2}$ loss of pairwise difference:
 $\mathcal{L}_{\textup{new}}(\phi){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{4}\,\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}{\mathbb{E}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;.$ (6)
+Therefore, to optimize the “real” reward model quality is to optimize an $\ell_{2}$ loss of pairwise difference:
+$-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}{\mathbb{E}}\Delta^{2}_{y,y^{\prime}}(r^{\star}-r_{\phi})\;.$ (7)
 Comparison with MLE.
 The reward model $r_{\phi}$ are typically learned via MLE from preference data, which does not consider the fact that the learned reward will ultimately be used to induce a policy.
+Let the distribution of the preference data be $\mu$ (by default $\mu$ is $\pi_{{\mathsf{ref}}}$, but can be any distribution here).
 Now revisit the MLE objective:
 $\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+\sigma(r^{\star}(y^{\prime})-r^{\star}(y))\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;,$
-whose gradient is:
+whose gradient is (see detailed calculations in Appendix B.12):
 $\nabla_{\phi}\mathcal{L}_{\textup{MLE}}(\phi)=-\underset{y,y^{\prime}\sim\mu}{\mathbb{E}}\left[\nabla_{\phi}r_{\phi}(y)-\nabla_{\phi}r_{\phi}(y^{\prime})\right]\left[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]\;.$ (8)
-Comparing Equation 4 with Equation 6, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 6 to further align with Equation 4.
+Following Equation 8, we can see that the gradient of DPO is
+$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\theta)\propto-\mathbb{E}_{y,y^{\prime}\sim\mu}[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))][\nabla(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))]\;,$
+and the gradient of reward modeling is
+$\nabla_{\phi}\mathcal{L}_{\text{RM}}(\phi)\propto-\mathbb{E}_{y,y^{\prime}\sim\mu}[\sigma(r^{\star}(y)-r^{\star}(y^{\prime}))-\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))][\nabla(r_{\phi}(y)-r_{\phi}(y^{\prime}))]\;.$
+Comparing Equation 5 with Equation 8, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 8 to further align it with Equation 5.
 And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$.
-And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$.
+And this by-product explains why is PILAF sampler (a variant of online sampler, see Definition 1) introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$.
 If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it can provide token-level reward information.
@@
 Since $r^{\star}\in\mathcal{F}$, RLHF exactly recovers $r^{\star}$ during reward learning.
-The policy optimization stage then solves $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi}$, so by definition, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$.
+The policy optimization stage then solves $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r^{\star}}^{\pi}$, so by definition, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}$.
 On the other hand, DPO is trained on preferences induced by $r^{\star}$.
@@
 B.3 Proof of Proposition 3
-Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}$.
+Since $r^{\star}\in\mathcal{F}$, RLHF recovers $r^{\star}$ exactly and then solves $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r^{\star}}^{\pi}$, by definition achieving $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}$.
 DPO, instead, minimizes a proxy loss defined over pairwise preferences.
-Since $\pi_{\textup{DPO}}\in\Pi$, we have $V_{r^{\star}}^{\pi_{\textup{DPO}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}$, which proves the first claim.
+Since $\pi_{\textup{DPO}}\in\Pi$, we have $V_{r^{\star}}^{\pi_{\textup{DPO}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V^{\Pi}_{r^{\star}}=V_{r^{\star}}^{\pi_{\textup{RLHF}}}$, which proves the first claim.
 For the strict gap, we consider a multi-armed bandit setting with the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
@@
 It then computes a policy $\pi_{\textup{RLHF}}$ that maximizes $V_{\hat{r}}^{\pi}$ over $\Pi$.
-In general, this policy is sub-optimal under $r^{\star}$ (see Proposition 3), and thus $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V_{\Pi}^{\star}$.
+In general, this policy is sub-optimal under $r^{\star}$ (see Proposition 3), and thus $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}=V^{\Pi}_{r^{\star}}$.
 DPO directly optimizes a preference-based loss over $\Pi$.
-Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V_{\Pi}^{\star}$.
+Since $\pi^{\star}\in\Pi$ and DPO is given access to exact preference data consistent with $r^{\star}$, it can recover $\pi^{\star}$, and hence $V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}=V^{\Pi}_{r^{\star}}$.
 For the strict gap, consider a multi-armed bandit setting analogous to Section B.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
@@
 This yields a strict gap:
-$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}\;.$
+$V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}\;.$
 B.5 Proof of Proposition 6
@@
 Since the exact solution for online DPO is hard to compute, we didn’t find elegant proofs for these two propositions. They are examined correct numerically.
-By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{\Pi}^{\star}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
+By Proposition 3, we have $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V^{\Pi}_{r^{\star}}=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$.
 Now we construct an environment under Condition 2, such that online DPO cannot outperform DPO, even with PILAF sampler.
@@
 Figure 5: Numerically Computed Curves of Gradient Functions and Value Functions.
-B.10 Formal Statement of Theorem 10 and Proof
-B.10.1 Preliminaries of single-token prediction
+B.10 Calculation of Token-level Structure of the Optimal Solution for DPO.
+As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
+$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,\;(t\neq 0)$
+$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\star}(y_{0})-\beta\log Z}{\beta}\right)\;,$
+where $Z:=\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$, and the $q^{\star}$ function is determined in a recursive way:
+[…]
+To prove this, we define a $q^{\prime}$ function as:
+$q^{\prime}(y_{0})=\beta\log Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;.$
+For the initial token, by definition we have:
+$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\prime}(y_{0})-\beta\log Z}{\beta}\right)\;.$ (10)
+And then for a $y$ with $y_{N}$ as the terminal token, we have:
+$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}$ $=\sum_{t=0}^{N}\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}$
+$=q^{\prime}(y_{0})-\beta\log Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
+$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
+Note that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, we have:
+$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=-\beta\log Z+r^{\star}(y)\;,$
+thus
+$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (11)
+Then by definition:
+$q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
+we have:
+$\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=\pi^{\star}(y_{t}|y_{0\ldots t-1})\;,$ (12)
+and thus
+$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(s|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
+which yields:
+$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s\in\mathcal{V}}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (13)
+Combining Equations 10, 12, 11 and 13, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
+Furthermore, if we assume $r^{\star}(y_{0\ldots t})$ to be decomposed as $\sum_{i=0}^{t}r^{\star}(y_{i}|y_{0\ldots i-1})$, we can define a soft value function $v^{\star}$:
+[…]
+Then samely, we have
+$\pi^{\star}(y_{t}|y_{0,\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0,\ldots t-1})\exp\left(\frac{r^{\star}(y_{t}|y_{0\ldots t-1})+v^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
+Proof is the same as what we do on $q^{\star}$ so we omit it.
+B.11 Formal Statement of Theorems 10 and 11 and Proofs
+B.11.1 Preliminaries of single-token prediction
 Before proceeding, we first prepare some ingredients for the single-token prediction task.
@@
 Proof of this lemma is given in Section B.11.4.
-B.10.2 Formal statement of Theorem 10
+B.11.2 Formal statement of Theorem 10
 Assumption 12 (Task configuration).
-Recall that in DTSP task, we have $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)$, where $\psi(y),\psi(y,\omega),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$.
-We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(y)\|_{2}\leq L$, and $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$.
+Recall that in DTSP task, we have $r^{\star}(a,b)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\beta e_{1}^{\top}\psi(a,b)$, where $a,b\in\mathcal{V}$, $\psi(a),\psi(a,b),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$.
+We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(a)\|_{2}\leq L$, and $\psi(a,b)=\psi(b)+(\textbf{r}_{\textup{dense}}^{\top}\psi(a))e_{1}$.
 Assumption 13 (Preference data collection).
-For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs.
-Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)}$ and $y_{l}^{(i)}=\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}$.
-And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))^{\top}$ is non-singular, and $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$.
+For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{a_{1}^{(i)},a_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{a_{1}^{(i)}a_{1}^{(i)},a_{2}^{(i)}a_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs.
+Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=a_{1}^{(i)}a_{1}^{(i)}$ and $y_{l}^{(i)}=a_{2}^{(i)}a_{2}^{(i)}$ w.p.
+$\sigma(r^{\star}(a_{1}^{(i)},a_{1}^{(i)})-r^{\star}(a_{2}^{(i)},a_{2}^{(i)}))$.
+And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))^{\top}$ is non-singular, $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$, and $n\geq c_{1}\text{tr}(\Sigma_{\mathcal{D}}^{-1})$, where $c_{1}$ is the constant in Lemma 1.
 Theorem 14 (Formal separation theorem).
-Under token-level linear parameterization and Assumptions 11 and 12, there exists an environment for DTSP task, s.t.
+Under token-level linear parameterization and Assumptions 12 and 13, there exists an environment for DTSP task, s.t.
 by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator:
@@
 $\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
-B.10.3 Proof of Theorem 13
-Let $\pi_{{\mathsf{ref}}}(\cdot|y)$ be identical for all $y$, then we have
-$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(y)+C_{5}\;,$
-for $\forall y$, where $C_{5}\in\mathbb{R}$ is an offset.
-Recall that:
-$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
-$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
+B.11.3 Proof of Theorem 14
+Let $\pi_{{\mathsf{ref}}}(\cdot|a)$ be identical for all $a$, then we have
+$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(\psi(a,b)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(a)+C_{5}\;,$
+for $\forall a\in\mathcal{V}$, where $C_{5}\in\mathbb{R}$ is a constant offset.
+Note that for the second token, $\theta_{r}^{\star}$ and $\theta_{p}^{\star}$ share the same optimal solution:
+$(\theta_{r,1}^{\star})^{\top}\psi(a,b)=e_{1}^{\top}\psi(a,b)+C_{1}\;,$
+$(\theta_{p,1}^{\star})^{\top}\psi(a,b)=e_{1}^{\top}\psi(a,b)+C_{2}\;,$
+where $C_{1},C_{2}\in\mathbb{R}$ are constant offsets when fixing $a$. And for the first token $a$, there is a distinction:
+$(\theta_{r,0}^{\star})^{\top}\psi(a)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+C_{3}\;,$
+$(\theta_{p,0}^{\star})^{\top}\psi(a)$ $=\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(r^{\star}(a,b)/\beta)+C_{4}$
+$=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(\psi(a,b)_{1})+C_{4}$
+$=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\textbf{r}_{\textup{dense}}^{\top}\psi(a)+C_{4}+C_{5}\;,$
+where $C_{3},C_{4}$ are constant offsets.
 We thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ and $\theta^{\star}_{p,0}=\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
 We can have a $\ell_{1}$-regularized estimator for the reward model:
-$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)})-r_{\theta}(\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
-$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
+$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(a_{w}^{(i)}a_{w}^{(i)})-r_{\theta}(a_{l}^{(i)}a_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
+$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
 where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 4 implies there exists appropriate $\gamma$, such that w.p. $1-\delta$,
-$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
+$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
 and thus w.p. $1-\delta$,
 $\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{w}^{(i)})-r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{l}^{(i)}))\right]^{2}$
-$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\theta_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
-$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\theta_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))-(\hat{\theta}_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
 $=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
 Note that
-$\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
-then Lemma 1 implies that $\exists N=\mathcal{O}(d)$, for $\forall n\geq N$ and any estimator $\hat{\theta}_{p}$, we have
-$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+$\log\sigma(\hat{r}_{\theta_{p}}(a_{w}^{(i)}a_{w}^{(i)})-\hat{r}_{\theta_{p}}(a_{l}^{(i)}a_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)})))\;,$
+then Lemma 1 implies that for any estimator $\hat{\theta}_{p}$, we have
+$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
 Now observe the data-induced semi-norm of surrogate reward learning:
-$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
-$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
-$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}\;.$
+$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(\hat{r}_{\hat{\theta}_{p}}(y_{w}^{(i)})-\hat{r}_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
+$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}\;.$
 And thus there exists an environment for DTSP, s.t.
-$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
-B.10.4 Proof of Lemma 4
+$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(\hat{r}_{\hat{\theta}_{p}}(y_{w}^{(i)})-\hat{r}_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
+B.11.4 Proof of Lemma 4
 Lemma 5 (Lemma D.4 of Yao et al. [2025]).
@@
 for $\forall\theta^{\prime}\in\mathbb{R}^{d}$ s.t. $\theta^{\prime}+\theta^{\star}\in\Theta_{B}$.
+Proof.
 We take $\gamma=\Theta\left(\sqrt{\frac{\log(d)+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ is determined in Theorem 3.3 of Yao et al.
@@
 $\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
-B.11 Omitted Calculations
-Calculation of token-level structure of the optimal solution for DPO. As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
-$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
-$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\star}(y_{0})-\beta\log Z}{\beta}\right)\;,$
-where $Z:=\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$, and the $q^{\star}$ function is determined in a recursive way:
-[…]
-To prove this, we define a $q^{\prime}$ function as:
-$q^{\prime}(y_{0})=\beta\log Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;.$
-For the initial token, by definition we have:
-$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\prime}(y_{0})-\beta\log Z}{\beta}\right)\;.$ (9)
-And then for a $y$ with $y_{N}$ as the terminal token, we have:
-$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}$ $=\sum_{t=0}^{N}\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}$
-$=q^{\prime}(y_{0})-\beta\log Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
-$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
-Note that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, we have:
-$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=-\beta\log Z+r^{\star}(y)\;,$
-thus
-$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (10)
-Then by definition:
-$q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
-we have:
-$\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=\pi^{\star}(y_{t}|y_{0\ldots t-1})\;,$ (11)
-and thus
-$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
-which yields:
-$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (12)
-Combining Equations 9, 11, 10 and 12, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
+B.11.5 Formal statement of Theorem 11 and proof
+Lemma 6 (Lemma J.5 of Nika et al. [2024]).
+If the features $\psi(a)$ are sampled from a 0-mean distribution and span $\mathbb{R}^{d}$, then $\log\sum_{a}\exp(\theta^{\top}\psi(a))$ is $\kappa$-strongly convex w.r.t.
+$\theta\in\Theta_{B}$, where $\kappa$ is an $\mathcal{O}(1)$ constant determined by $\beta,B,L$ and $|\mathcal{V}|$.
+Theorem 15 (Formal sub-optimality separation theorem).
+In the same setting as Theorem 14, there exists an environment for DTSP task, s.t.
+the sub-optimality of the RLHF policy model $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\hat{\theta}_{r}}}^{\pi}$ can be reduced to $\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)$, i.e.
+w.p.
+$1-\delta$,
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $=\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)\;,$
+while the sub-optimality of the DPO policy model $\pi_{\textup{DPO}}=\pi_{\hat{\theta}_{p}}$ is lower bounded:
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $=\Omega\left(\frac{d}{n}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}\right)\;.$
+Proof.
+The proof follows the ideas of Theorem 3.2 of Zhu et al. [2023] and Theorem 4.2 of Nika et al. [2024], with appropriate adaptations to our setting.
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $\leq\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},b_{1}\sim\pi^{\star}(\cdot|a_{1}),\\ a_{2}\sim\pi_{\textup{RLHF}},b_{2}\sim\pi_{\textup{RLHF}}(\cdot|a_{2})\end{subarray}}{\mathbb{E}}\left[\left(r^{\star}(a_{1},b_{1})-r^{\star}(a_{2},b_{2})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(a_{1},b_{1})}{\pi_{{\mathsf{ref}}}(a_{1},b_{1})}-\beta\log\frac{\pi_{\text{RLHF}}(a_{2},b_{2})}{\pi_{{\mathsf{ref}}}(a_{2},b_{2})}\right)\right]$
+$=\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},b_{1}\sim\pi^{\star}(\cdot|a_{1}),\\ a_{2}\sim\pi_{\textup{RLHF}},b_{2}\sim\pi_{\textup{RLHF}}(\cdot|a_{2})\end{subarray}}{\mathbb{E}}\left[\left(r^{\star}(a_{1},b_{1})-r^{\star}(a_{2},b_{2})\right)-\left(r_{\hat{\theta}_{r}}(a_{1},b_{1})-r_{\hat{\theta}_{r}}(a_{2},b_{2})\right)\right]$
+$=\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}\left[\beta(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})^{\top}(\psi(a_{1})-\psi(a_{2}))\right]$
+$=\beta(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})^{\top}\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}(\psi(a_{1})-\psi(a_{2}))$
+$\leq\beta\|\Sigma_{\mathcal{D}}^{1/2}(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})\|_{2}\|\Sigma_{\mathcal{D}}^{-1/2}\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}(\psi(a_{1})-\psi(a_{2}))\|_{2}$
+$=\beta\|\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0}\|_{\Sigma_{\mathcal{D}}}\cdot\mathcal{O}\left(\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)$
+$=\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)\;.$
+The first inequality comes from performance difference lemma (see Section B.12); the second equality comes from the observation that all $r_{\theta_{r}}$ with $\theta_{r,1}=e_{1}$ can be fitted by the log-linear policy model; the third and fourth equalities come from simple calculations under our setting; the fifth inequality comes from Cauchy-Schwarz inequality; the sixth equality comes from the fact that $\psi(a)$ is bounded; and the last equation comes from Theorem 14.
+Since the optimal policy satisfies $\pi^{\star}(a,b)=\pi_{{\mathsf{ref}}}(a,b)\exp(r(a,b)/\beta)/Z$, we have:
+$V_{r^{\star}}^{\pi^{\star}}$ $=\underset{a\sim\pi^{\star},b\sim\pi^{\star}(\cdot|a)}{\mathbb{E}}\left[r^{\star}(a,b)-\beta\log\frac{\pi^{\star}(a,b)}{\pi_{{\mathsf{ref}}}(a,b)}\right]$
+$=\beta\log Z$
+$=r^{\star}(a^{\prime},b^{\prime})-\beta\log\frac{\pi^{\star}(a^{\prime},b^{\prime})}{\pi_{{\mathsf{ref}}}(a^{\prime},b^{\prime})}\;,\;\forall a^{\prime},b^{\prime}\in\mathcal{V}\;.$
+Then we have:
+$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $=\underset{a\sim\pi_{\textup{DPO}},b\sim\pi_{\textup{DPO}}(\cdot|a)}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(a,b)}{\pi_{{\mathsf{ref}}}(a,b)}-r^{\star}(a,b)+V_{r^{\star}}^{\pi^{\star}}\right]$
+$=\underset{a\sim\pi_{\textup{DPO}},b\sim\pi_{\textup{DPO}}(\cdot|a)}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(a,b)-\beta\log\pi^{\star}(a,b)\right]$
+$=\beta\underset{a\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\left(\hat{\theta}_{p,0}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right]+\beta\log\frac{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}\;,$
+where $v$ can be any vector in $\mathbb{R}^{d}$.
+Recall that we require $\pi_{{\mathsf{ref}}}(\cdot|a)$ to be identical for all $a\in\mathcal{V}$ in the proof of Theorem 14.
+Here we further construct $\pi_{{\mathsf{ref}}}$ to be uniform on the first token.
+Now observe
+$\log\frac{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}$ $=\log\frac{\sum_{a\in\mathcal{V}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\sum_{a\in\mathcal{V}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}\;.$
+Set $v$ to be $\frac{1}{|\mathcal{V}|}\sum_{a\in\mathcal{V}}\psi(a)$, then we have $\sum_{a\in\mathcal{V}}\left(\psi(a)-v\right)=0$.
+Since $\Sigma_{\mathcal{D}}$ is already non-singular, we have that $\{\psi(a)-v\}_{a\in\mathcal{V}}$ can span $\mathbb{R}^{d}$.
+So we can directly apply Lemma 6, and get
+$\log\sum_{a\in\mathcal{V}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)-\log\sum_{a\in\mathcal{V}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)$
+$\geq$ $\langle\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0},\nabla_{\theta}\log\sum_{a\in\mathcal{V}}\exp\left(\theta^{\top}\left(\psi(a)-v\right)\right)|_{\theta=\hat{\theta}_{p,0}}\rangle+\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}$
+$=$ $-\underset{a\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\left(\hat{\theta}_{p,0}-(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}})\right)^{\top}\left(\psi(a)-v\right)\right]+\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}\;.$
+Therefore, we have
+$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $\geq\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}$
+$=\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}\|\Sigma_{\mathcal{D}}\|_{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
+$\geq\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}\left\|\Sigma_{\mathcal{D}}\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right)\right\|_{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
+$\geq\frac{\kappa}{2}\langle\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0},\Sigma_{\mathcal{D}}\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right)\rangle\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
+$=\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{\Sigma_{\mathcal{D}}}^{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
+$=\Omega\left(\frac{d}{n}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}\right)\;.$
+The first inequality comes from Lemma 6; the second equality comes from the non-singularity of $\Sigma_{\mathcal{D}}$; the third inequality comes from a standard property of the spectral norm; the fourth inequality comes from Cauchy-Schwartz inequality; the fifth equality is a simple algebraic equality; and the last equation comes from Theorem 14.
+B.12 Omitted Calculations
+Calculation of the sub-optimality with respect to the mis-specification error.
+First, note that for any $\pi\in\Delta(\mathcal{Y})$, we have:
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi}$ $=\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}\right]-\underset{y\sim\pi}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]\;,$
+$=\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi^{\star}(y)}{\pi(y)}-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]-\underset{y\sim\pi}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]$
+$=-\mathsf{KL}\left(\pi^{\star}\|\pi\right)+\underset{y\sim\pi^{\star},y^{\prime}\sim\pi}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
+$\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]\;.$
+We call it the performance difference lemma (Lemma 1 of Shi et al. [2025]).
+For RLHF, we have:
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{RLHF}}}$ $\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi_{\text{RLHF}}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
+$\leq\max_{y,y^{\prime}\in\mathcal{Y}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
+$\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]}_{\text{reward model mis-specification error}}$
+$\quad+\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))-\left(\beta\log\frac{\pi_{\text{RLHF}}(y|x)}{\pi_{{\mathsf{ref}}}(y|x)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]}_{\text{policy model mis-specification error}}\;,$
+where the first inequality is by performance difference lemma, and the last two inequalities are by symmetry and the properties of $\max$.
+And if $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, by the definition of $\pi_{\textup{RLHF}}$, we have
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{RLHF}}}\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]}_{\text{reward model mis-specification error}}\;.$
+For DPO, by performance difference lemma, we have:
+$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{DPO}}}$ $\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi_{\text{DPO}}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{DPO}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
+$\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{DPO}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]}_{\text{policy model mis-specification error}}$
+$=\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\text{DPO}}(y)-\hat{r}_{\text{DPO}}(y^{\prime}))\right]}_{\text{surrogate reward model mis-specification error}}\;.$
+The first inequality is by performance difference lemma, the second inequality is by symmetry and the property of $\max$, and the last equality is just another interpretation.
+Therefore, we can see that the sub-optimality of each algorithm can be upper bounded by the linear model mis-specification error.
 Calculation of the underlying “real” objective.
@@
 $\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
-Below we provide two additional results for reference.
-Exact Optimization: calculation of the policy performance gap. Since the optimal policy satisfies $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
-$V_{r^{\star}}^{\pi^{\star}}=\underset{y\sim\pi^{\star}}{\mathbb{E}}r(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=\beta\log Z=r^{\star}(y^{\prime})-\beta\log\frac{\pi^{\star}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\;,\;\forall y^{\prime}\in\mathcal{Y}\;.$
-Then as motivated by Nika et al. [2024], we can split the performance gap of two-stage RLHF as:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
-and for DPO we also have:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
-Approximate Optimization: calculation of the reward representation gap.
-Motivated by Nika et al.
-[2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi^{\star}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
-$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$ ($\mathcal{F}\subseteq\mathcal{F}_{\Pi}$)
-$=\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
-and for DPO we have:
-$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi^{\star}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
-$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
-$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
 Appendix C Implementation Details
@@
 Our reward model is GPT2-large-harmless model [Yang et al., 2024] (https://huggingface.co/Ray2333/gpt2-large-harmless-reward_model).
+More implementation details.
+For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset.
+For approximate optimization, we instead compute the empirical BT loss.
+We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$.
+During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$.
+Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
 Hyper-parameters.